### v1

# Feature Fusion from Head to Tail: an Extreme Augmenting Strategy for Long-Tailed Visual Recognition

Abstract
The imbalanced distribution of long-tailed data poses a challenge for deep neural networks, as models tend to prioritize correctly classifying head classes over others so that perform poorly on tail classes. The lack of semantics for tail classes is one of the key factors contributing to their low recognition accuracy. To rectify this issue, we propose to augment tail classes by borrowing the diverse semantic information from head classes, referred to as head-to-tail fusion (H2T). We randomly replace a portion of the feature maps of the tail class with those of the head class. The fused feature map can effectively enhance the diversity of tail classes by incorporating features from head classes that are relevant to them. The proposed method is easy to implement due to its additive fusion module, making it highly compatible with existing long-tail recognition methods for further performance boosting. Extensive experiments on various long-tailed benchmarks demonstrate the effectiveness of the proposed H2T. The source code is temporarily available at https://github.com/Keke921/H2T.
1 Introduction
(a) Comparison of class boundaries. The orange line represents the decision boundary. $\bullet$ and $\blacktriangle$ denotes head and tail class samples, respectively.
(b) Frequency of predictive labels for tail classes samples. A large number of tail class samples are incorrectly recognized as head classes.
Figure 1: The relationship between tail and head classes.
Deep neural networks (DNNs) have shown remarkable capabilities in diverse visual recognition tasks [16, 23, 6, 60, 34], yet their performance heavily relies on data that evenly distribute across categories. In contrast, real-world data typically exhibit long-tailed distributions [37, 33, 54], which has become one of the bottlenecks limiting the advancement of DNNs and hinders their broader practical applications. For example, misdiagnosis of rare diseases by computer-aided diagnosis [45, 2] may result in irreversible consequences in severe cases. Recently, a large number of methods have been proposed to address the class imbalance problem in long-tailed data. Class-balancing methods, including re-sampling [5, 18, 39, 20, 35] and re-weighting [17, 12, 4, 30, 15], are the most straightforward approaches to tackle class imbalance. Previous class-balancing methods, however, improve the overall performance but are usually at the risk of overfitting to the minority classes or the loss of information [52, 36, 20]. To overcome these limitations, recent studies have proposed alternative techniques such as re-margining methods [4, 27, 26] and ensembling learning [47, 24, 57, 10]. These methods can generate more robust predictions, but they do not consider enhancing the sample diversity of tail classes. However, insufficient samples and a lack of tail class semantics will also lead to a low recognition rate for tail classes. Data augmentation methods [46, 36, 20, 7, 28] are thus proposed to effectively enrich tail classes. While these methods have shown state-of-the-art performance, they generally aim to obtain a more representative and separable feature space and do not dig into the potential of the classifier in the obtained feature space.
This paper thereby aims to unleash the power of the obtained feature space to get a better classifier. Take the binary case in the obtained feature space as an example, the semantic space can be divided into three areas: head-like, tail-like, and hard negative. The former two areas represent the main distribution spaces of the head and tail classes, respectively, while the hard negative area includes samples that are difficult to distinguish between the two classes. In long-tail data, the head class is usually sufficiently sampled, and its semantic area is fully occupied. Besides, some samples of head class are distributed in the hard negative area. However, only a few samples are often distributed in the tail class semantic regions due to insufficient tail class sampling, resulting in a decision boundary that is limited to the semantic area of the tail class, as shown in the upper of Figure 1a.
To address the issue of the lack of tail class semantics, we propose a direct and effective solution, named head-to-tail fusion (H2T), which borrows partial of the semantics from the head class and applies them to the tail class. Our approach is based on the assumption that predictions on rare instances are easily affected by objects that appear frequently. Figure 1b shows a simple experiment [^1] The experiment was conducted on CIFAR10-LT with an imbalance ratio of 100. The backbone network is ResNet-32 which was trained using cross-entropy loss. Classes 7-9 (0-indexed) are tail classes.for this assumption. The label frequency of predicting tail classes into each class is presented. We can observe that a large number of tail classes are misclassified as head classes by the model, which may indicate that the hard negative area for tail classes is filled with head class samples. Therefore, there is relatively more common closet semantics between head and tail classes. By transferring such semantics, we can effectively fill the tail semantic area and improve the recognition rate of the tail class samples, as shown in the lower of Figure 1a. On the other hand, by obtaining partial specific semantics (not common semantics) of the head class, the tail class samples can occupy the hard negative semantic region, which forces the classification boundary to move from the tail semantic region to the hard negative semantic region. This approach brings the classification boundary closer to the theoretical decision boundary, and improves the classification performance for tail classes. To streamline the fusion operation, we design a simple and easy-to-implement strategy that takes full use of the obtained features without retraining the backbone. Specifically, we randomly substitute several channels of the feature maps of the class-balanced data with that of the head-biased data. This makes the feature maps of tail classes with a high probability of being fused with head classes. H2T can be flexibly integrated with other methods, as it does not alter the structure of the backbone as well as any other technical details. The main contributions of this paper are summarized as follows:
  • We propose a novel H2T that borrows information from head classes and applies it to augment tail classes without requiring additional data or network parameters, which can enrich tail classes and release the potential of obtained feature space.
  • We devise a simple fusion strategy that partially substitutes the feature maps of an instance-wise sampled data branch with those of a class-balanced sampled branch. The adoption of this approach enables H2T to be implemented with ease and without requiring any modifications to the original model structure.
  • Considerable performance improvement is achieved in both single and multi-expert models by integrating the H2T module into the original structure. Extensive experiments on popular benchmarks validate the efficacy of the proposed method.
2 Related Work
This section mainly makes an overview of methods based on data augmentation that is the most relevant regime.
Classical Augmentation: Classical augmentation methods including flip, rotate, crop, padding, and color jittering etc. [41, 13] have been widely applied in the process of deep model training. On long-tailed data, these methods can improve model robustness and prevent overfitting to specific input patterns to a certain extent. Recently, AutoAugment [8], Fast AutoAugment [29], Population Based Augmentation [14] and Randaugment [9], etc. have been proposed to automatically determine the best data augmentation strategy for a dataset. These methods have been shown to increase classification accuracy for DNNs and demonstrated their efficacy on long-tailed data [38, 11, 52, 25].
MixUp Based Augmentation: MixUp [53] is an effective data augmentation method that linearly combines a pair of row images and their labels. This method adds complexity control to the uncovered space in the data space through linearly interpolating discrete sample points, which reduces the model generalization error. Numerous studies [55, 56, 27] have experimentally proven that MixUp can significantly improve long-tail recognition performance. CutMix [50] is a variant of MixUp that augments data by cutting out a patch from the input and pasting it onto another image in the training set. The ground truth labels are combined in proportion to the patch areas. Park et al. [36] proposes to mix the foreground patches from tail classes with the background images from head classes via CutMix, enabling the transfer of context-rich background information from head to tail.
New Images Based Augmentation: In addition to the aforementioned methods, researchers have proposed alternative approaches that generate new informative samples for tail classes. One such method is M2m [20], which translates samples from head classes to the tail using a pre-trained auxiliary classifier. Wang et al. [46] utilize the learned encoded variation information from head classes to synthesize images for tail classes. Zada et al. [52] directly exploit pure noise images as tail class samples. Zhang et al. [55] utilize class activation maps (CAM) [58] to separate image foreground and background, followed by augmenting the foreground through flipping, rotating and scaling, etc. The new samples are then generated by overlaying the augmented foreground on the unchanged background.
Feature Level Augmentation: Data augmentation can also be performed in feature space. For example, Manifold MixUp [44] conducts the linear interpolation on the features of input images, which has been shown to yield better results in long-tail learning [55, 57]. Chu et al. [7] use CAM to decompose features into class-generic and class-specific components. Tail classes are then augmented by combining their class-specific components with the class-generic components. Another promising technique is to transfer knowledge from the head classes to the tail classes. For example, Yin et al. [49] and Liu et al. [32] enrich tail classes with the intra-class variance of head classes to balance the class distributions in feature space. Recently, meta-learning-based augmentation has been leveraged to address class imbalance. Liu et al. [31] design a meta-embedding that uses a memory bank to enrich tail classes. MetaSAug [28] learns class-wise covariance matrices by minimizing the loss on a small balanced validation set and subsequently uses the learned matrices to semantically augment tail classes.
3 Methodology
3.1 Preliminaries
We formally define the basic notations used in this paper before going into detail about our proposed method.
For dataset: We use $\{x,y\}$ to represent one input image and its corresponding label. The total number of classes is denoted by $C$, and the training set includes $N$ samples. Thus, we have $y\in\{0,1,\cdots,C-1\}$. Suppose class $i$ has $n_{i}$ training samples. Then $N=\sum_{i}n_{i}$. For simplicity, we suppose $n_{0}\geq n_{1}\geq\cdots\geq n_{C-1}$
For backbone network: Feed $x$ into the DNN backbone, we can obtain its feature maps before the last pooling layer, denoted as $\mathcal{F}=\left[F_{1},F_{2},\cdots,F_{d}\right]\in\mathbb{R}^{w_{F}\times h_{F}\times d}$, where $w_{F}$ and $h_{F}$ represent the width and height of the feature map, respectively. $d$ is the dimension of features in embedding space. The representation after the last pooling layer is ${f}\in\mathbb{R}^{d}$. The weight of the linear classifier is represented as $\textbf{W}=[w_{1},w_{2},\cdots,w_{c}]\in\mathbb{R}^{d\times C}$, where $w_{i}$ represents the classifier weight for class $i$. We use the subscripts $h,m$ and $t$ to indicate head, medium and tail classes, respectively. $z_{i}=w_{i}^{T}f$ represents the predicted logit of class $i$, where the subscript $i=y$ denotes the target logit and $i\neq y$ denotes the non-target logit.
3.2 Motivation
Figure 2: Framework of head-to-tail fusion (H2T).
We begin by analyzing the compression of the tail distribution caused by imbalances in class sizes. The standard training method for DNNs typically involves either empirical risk minimization (ERM) [43, 51] or structural risk minimization (SRM) [22]. Since the extra item in SRM is a function of model parameters, we can analyze the effect of data on DNNs without loss of generality by focusing solely on ERM. To facilitate this analysis, we partition the data by class and use the superscript $i$ to indicate their corresponding labels. Formally, ERM can be expressed as:
$\mathcal{R}_{emp}(\phi)=\frac{1}{N}\sum_{i=1}^{C}\sum_{j=1}^{n_{i}}\mathcal{L}(\phi(x^{i}_{j}),i)$ (1)
where $\mathcal{L}$ is the loss function. $n_{h}\gg n_{t}$ where $h<t$ Thus, misclassifying tail classes have much less risk compared with head classes. Generally, a considerable number of tail class samples are mislabeled as head classes. The straightforward solution involves increasing the importance or frequencies of tail classes. Re-weighting (RW) [4] or re-sampling (RS) [57, 1] can increase the risk of misclassification for tail classes by assigning larger weights or higher sampling rates to them. This, in turn, increases the risk of overfitting to tail classes. To expand the spatial distribution of tail classes and prevent overfitting, it is crucial to enhance the diversity of tail class samples. Unfortunately, obtaining more samples is generally infeasible. And Figure 1b shows that the tail classes are more prone to confusing with head classes. Typically, there are multiple instances of head and tail classes that are intermixed around the decision boundary. We leverage these intermixed instances by directly borrowing semantic information from head classes to augment tail classes. We have two purposes: i) increase the risk of misclassifying tail class samples to prevent the model from simply labeling them as head classes; ii) transfer the rich semantic information from head to tail classes, thereby improving the diversity of tail classes.
3.3 Fusing Head Features to Tail
Due to the imbalanced distribution of class samples, with the head classes being significantly more numerous than the tail classes, the latter tend to be crowded and underrepresented, leading to poor generalization performance on the test set of tail classes. As a large amount of tail class samples are misclassified as head classes, there may exist closet semantics between head and tail classes. To exploit this abundant closet semantic information, we fuse the features of head classes to the tail. This operation achieves the purpose of enriching the tail classes and expanding their feature space distribution. The fusion process is formulated as:
~$\tilde{\mathcal{F}}=\mathcal{M}_{p}\otimes\mathcal{F}^{t}+(\mathbf{I}-\mathcal{M}_{p})\otimes\mathcal{F}^{h},$ (2)
where $\mathcal{M}_{p}$ is the mask combined with identity matrices $\mathbf{I}_{m}\in\mathcal{R}^{w_{F}\times h_{F}}$ and zero matrices $\mathbf{O}_{m}\in\mathcal{R}^{w_{F}\times h_{F}}$. The parameter $p$ represents the fusion ratio. Specially, the number of $\mathbf{I}_{m}$ and $\mathbf{O}_{m}$ is $d\cdot p$ and $d\cdot(1-p)$, respectively. $\tilde{\mathcal{F}}$ is then passed through a pooling layer and classifier to predict the corresponding logits $\mathbf{z}=[z_{1},z_{2},\cdots,z_{C}]$. Different loss functions such as CE loss, MisLAS [56], or GCL [27], to name a few, can be adopted. The backbone $\phi$ can choose the single model [13] as well as multi-expert model [47, 48]. We exploit the two-stage method and apply H2T in stage II.
Input: Training set, fusion ration $p$ ;
Output: Trained model;
1 Initialize the model $\phi$ randomly ;
2 for $iter=1$ to $E_{0}$ do
3       Sampling batches of data $(x,y)\sim\mathcal{T}^{I}$ from the instance-wise sampling data;
4       Obtain the feature map $\mathcal{F}=\phi_{\theta}(x)$ ;
5       Calculate logits $z=\mathbf{W}^{T}\mathbf{f}$ and loss $\mathcal{L}_{1}(x,y)$;
6       $\phi=\phi-\alpha\nabla_{\phi}\mathcal{L}_{1}((x,y);\phi)$.
7 end for
8for $iter=E_{0}+1$ to $E_{2}$ do
9       Sample batches of data $(x^{B},y^{B})\sim\mathcal{T}^{B}$ and $(x,y)\sim\mathcal{T}^{I}$;
10       Obtain feature maps $\mathcal{F}^{B}=\phi_{\theta}(x^{B})$ and $\mathcal{F}^{I}=\phi_{\theta}(x)$ ;
11       Fuse feature maps by Eq. 2 to obtain $\tilde{\mathcal{F}}$ ;
12       Input $\tilde{\mathcal{F}}$ to pooling layer to obtain $\tilde{\mathbf{f}}$, then calculate the logits by $\tilde{z}=\mathbf{W}^{T}\tilde{\mathbf{f}}$ and the loss by $\mathcal{L}_{2}(x^{B},y^{B})$ ;
13       Froze the parameters of representation learning $\phi^{r}$, and finetune the classifier parameters $\phi^{c}$: $\phi^{c}=\phi^{c}-\alpha\nabla_{\phi^{c}}\mathcal{L}_{2}((x^{B},y^{B});\phi^{c})$.
14 end for
Algorithm 1 H2T
Visual recognition tasks often involve a vast number of categories. Manually screening fused categories can be a tedious and time-consuming process. We thereby design a simple strategy. We sample two versions of data: 1) class-balanced data $\mathcal{T}^{B}$ used to balance the ERM of each class, which is fed into the fused branch, and 2) instance-wise data $\mathcal{T}^{I}$ has a high probability to obtain head class samples, which is fed into the fusing branch. The sampling rates $p^{B}_{i}$ and $p^{I}_{i}$ for $\mathcal{T}^{B}$ and $\mathcal{T}^{I}$ are calculated by
$p^{B}_{i}=\frac{1}{C\cdot n_{i}},\>p^{I}_{i}=\frac{1}{N}.$ (3)
Balanced sampling data ensures that each class is sampled with equal probability $\frac{1}{C}$. The fewer the number of samples, the higher the probability of being resampled multiple times. The original data is distributed with a long tail, resulting in a head-biased dataset where the head classes have a higher probability of being sampled. Next, we can use the feature maps $\mathcal{F}^{B}$ and $\mathcal{F}^{I}$ obtained from $\mathcal{T}^{B}$ and $\mathcal{T}^{I}$ to substitute $\mathcal{F}^{t}$ and $\mathcal{F}^{h}$ in Eq. (2). By doing so, the features of the repeatedly sampled tail classes will be fused with the head class features with a higher probability. Different from the common practice of linearly combining a pair of inputs and their labels for augmentation, we only use the labels of the fused branch, namely the balance-sampled data, as ground truth to utilize the semantics of head classes to enrich tail classes. The proposed framework is illustrated in Figure 2 and Algorithm 1. The key steps are highlighted in rose-red arrow and text, respectively.
3.4 Rationale Analysis
We explore the theoretical rationality of the seemingly intuitive operation of H2T in depth in this section. For ease of analysis, we assume that the feature maps to be fused are rearranged in order without loss of generality. Then, after the pooling layer, the feature can be written as $f_{i}^{T}=[\dot{f}_{i}^{T},\ddot{f}_{i}^{T}]$, where $\dot{f}_{i}$ and $\ddot{f}_{i}$ denote the portions of features that are retained and to be fused, respectively $w_{i}^{T}=[\dot{w}_{i}^{T},\ddot{w}_{i}^{T}]$ is the corresponding classifier weights.
On the one hand, we expect $z_{t}>z_{h}$ for tail class, so that
$\begin{split}w_{t}^{T}f_{t}&>w_{h}^{T}f_{t}\>\>\Rightarrow\\ \dot{w}_{t}^{T}\dot{f}_{t}+\ddot{w}_{t}^{T}\ddot{f}_{t}&>\dot{w}_{h}^{T}\dot{f}_{t}+\ddot{w}_{h}^{T}\ddot{f}_{t}.\end{split}$ (4)
However, as shown in Figure 1b and analyzed in Section 3.2, numerous tail class samples are incorrectly classified into head classes. Therefore, the trained model actually predicts $z_{h}>z_{t}$, namely,
$\dot{w}_{t}^{T}\dot{f}_{t}+\ddot{w}_{t}^{T}\ddot{f}_{t}<\dot{w}_{h}^{T}\dot{f}_{t}+\ddot{w}_{h}^{T}\ddot{f}_{t}.$ (5)
After fusing the head to the tail in stage II, the feature of tail class is $\tilde{f}_{t}=[\dot{f}_{t},\ddot{f}_{h}]$, and the corresponding logit is $\tilde{z}_{t}$. Our training goal is still to make the target logit larger than the non-target class, namely, $\tilde{z}_{t}>\tilde{z}_{h}$. Therefore, we have
$\begin{split}w_{t}^{T}\tilde{f}_{t}&>w_{h}^{T}\tilde{f}_{t}\Rightarrow\\ \begin{bmatrix}\dot{w}_{t}\\ \ddot{w}_{t}\end{bmatrix}^{T}\begin{bmatrix}\dot{f}_{t}\\ \ddot{f}_{h}\end{bmatrix}&>\begin{bmatrix}\dot{w}_{h}\\ \ddot{w}_{h}\end{bmatrix}^{T}\begin{bmatrix}\dot{f}_{t}\\ \ddot{f}_{h}\end{bmatrix}\Rightarrow\\ \dot{w}_{t}^{T}\dot{f}_{t}+\ddot{w}_{t}^{T}\ddot{f}_{h}&>\dot{w}_{h}^{T}\dot{f}_{t}+\ddot{w}_{h}^{T}\ddot{f}_{h}.\end{split}$ (6)
Adding the last line in Eq. (6) to Eq. (5), we can obtain that
$\ddot{w}_{h}^{T}(\ddot{f}_{t}-\ddot{f}_{h})>\ddot{w}_{t}^{T}(\ddot{f}_{t}-\ddot{f}_{h}).$ (7)
We use $\ddot{\theta}_{*}$ to represent the angle between $\ddot{w}_{*}$ and $\ddot{f}_{t}-\ddot{f}_{h}$. Eq. (7) is further simplified to
$|\ddot{w}_{h}|\cos\ddot{\theta}_{h}>|\ddot{w}_{t}|\cos\ddot{\theta}_{t}.$ (8)
On the other hand, similar to Eq. (4), for logits of head classes, we have
$\dot{w}_{h}^{T}\dot{f}_{h}+\ddot{w}_{h}^{T}\ddot{f}_{h}>\dot{w}_{t}^{T}\dot{f}_{h}+\ddot{w}_{t}^{T}\ddot{f}_{h}.$ (9)
Add the last line in Eq. (6) to Eq. (9), we obtain that
$|\dot{w}_{t}|\cos\dot{\theta}_{t}>|\dot{w}_{h}|\cos\dot{\theta}_{h}.$ (10)
Eq. (8) forces $w_{h}$ to move closer to tail class samples while pushing $w_{t}$ further away from them, thus increasing the distribution span of head classes. Conversely, Eq. (10) exerts an opposite force. Figure 3 geometrically interprets the rationale of H2T. Notably, when the fusion ratio $p$ is small, Eq. (10) exerts a greater force, enabling the classifier to be well-calibrated. As $p$ grows, the item $\ddot{w}_{t}^{T}\ddot{f}_{h}>\ddot{w}_{h}^{T}\ddot{f}_{h}$ in Eq. (7) progressively becomes more dominant. Under this case, even though Eq. (8) encourages $w_{t}$ to move away from tail class samples, which is larger than the force of Eq. (10), $\ddot{w}_{t}^{T}\ddot{f}_{h}>\ddot{w}_{h}^{T}\ddot{f}_{h}$ attracts $w_{t}$ towards head class samples, thus broadens the distribution span of tail classes. Consequently, the classifier increases its performance in the tail class, regardless of the value of $p$. Section 4.4 also experimentally validates this analysis.
Figure 3: Rationale analysis of H2T. $\bullet$: $x_{h}$. $\blacktriangle$: $x_{t}$. $\rightarrow$: $w_{h}$. $\rightarrow$: $w_{t}$. $\rightarrow$: $f_{t}-f_{h}$. Forces 1⃝ and 2⃝ are generated by Eq. (10) and Eq. (8), respectively. 1⃝ $>$ 2⃝ makes the tail class sample to“pull" closer to $w_{t}$ and “push" further away from $w_{h}$. This leads to the enlargement of the tail class space and a reduction in the priority of head class.
4 Experiments
Table 1: Comparison results on CIFAR10/100-LT w.r.t. top-1 accuracy (%).
Dataset CIFAR10-LT CIFAR100-LT
Backbone: ResNet-32 Imbalance Ratio
Method Year 200 100 50 200 100 50
Single Model
CE loss - 67.24 70.70 74.81 35.99 39.55 45.40
LDAM-DRW [4] 2019 73.52 77.03 81.03 38.91 42.04 47.62
DR+MU [19] 2020 73.06 79.15 84.21 41.73 45.68 50.86
MisLAS [56] 2021 77.17 82.12 85.55 43.45 46.71 52.31
${}^{\star}$ [31] 2022 - 81.00 86.60 - 45.80 52.60
GCL [27] 2022 78.56 82.06 85.36 44.76 48.61 53.55
${}^{\star}$ [36] 2022 - - - - 43.90 48.30
BSCE [${}^{\star}$ [36] 2022 - - - - 46.60 51.40
[…]
MisLAS+H2T - 78.05 ($\uparrow$ 0.88) 83.23 ($\uparrow$ 1.11) 86.72 ($\uparrow$ 1.17) 43.84 ($\uparrow$ 0.39) 47.62 ($\uparrow$ 0.91) 52.73 ($\uparrow$ 0.42)
GCL+H2T - 79.42 ($\uparrow$ 0.86) 82.40 ($\uparrow$ 0.34) 85.41 ($\uparrow$ 0.05) 45.24 ($\uparrow$ 0.48) 48.88 ($\uparrow$ 0.27) 53.76 ($\uparrow$ 0.21)
Multi-Expert Model
BBN [57] 2020 73.47 79.82 81.18 37.21 42.56 47.02
RIDE [47] 2021 - - - 45.84 50.37 54.99
ACE [${}^{\star}$ 2021 - 81.20 84.30 - 49.40 50.70
ResLT [10] 2022 75.29 80.58 84.07 44.26 48.73 53.81
[…]
ResLT+H2T - 77.25 ($\uparrow$ 1.96) 81.77 ($\uparrow$ 1.19) 84.99 ($\uparrow$ 0.92) 46.18 ($\uparrow$ 1.92) 49.60 ($\uparrow$ 0.87) 54.39 ($\uparrow$ 0.58)
  • Note: The results are obtained by re-implementing the official codes or reproducing with the given parameters, except the methods marked by $\star$ that are quoted from the corresponding papers.
  • The best and the second-best results are shown in underline bold and bold, respectively.
4.1 Datasets and Metrics
Dataset: We evaluated our proposed method on five widely-used benchmarks: CIFAR10/100-LT [4], ImageNet-LT [33], iNaturalist 2018 [42], and Places-LT [33]. CIFAR10/100-LT are the small-scale datasets that are sub-sampled from their balanced versions, namely, CIFAR10/100 [21]. We used the imbalance ratios ($\rho=\frac{n_{max}}{n_{min}}$) 200, 100, and 50 [7]. The original versions of imageNet-2012 [40] and Places 365 [59] are also balanced datasets. We use the same settings as Liu et al. [33] to obtain the long-tailed versions. iNaturalist is a natural dataset collected from all over the world and is heavily imbalanced. The 2018 version [42] is utilized in our experiment.
Metrics: Besides top-1 classification accuracy, following Liu et al. [33], the accuracy on three partitions: head ($n_{i}>100$), medium ($20<n_{i}\leq 100$) and tail ($n_{i}\leq 20$) on large-scale datasets are also compared.
4.2 Basic Settings
SGD with a momentum of 0.9 is adopted for all datasets. Stage II is trained with 10 epochs. For CIFAR10/100-LT, we refer to the settings in Cao et al. [4] and Zhong et al. [56]. The backbone network is ResNet-32 [13]. Stage I trains for 200 epochs.The initial learning rate ($l_{r}$) is 0.1 and the batch size is 128. $l_{r}$ is decayed at the $160^{th}$ and $180^{th}$ epochs by 0.1. For imageNet-LT and iNaturalist 2018, we use the commonly used ResNet-50, while for Places-LT, we utilize the ResNet-152 pre-trained on imageNet. For other settings, we mainly follow the original papers of the comparison methods. We reproduce the prior methods that provide the hyper-parameters for a fair comparison. For those methods that do not provide hyperparameters or code, we directly report their results presented in their original paper.
4.3 Comparisons to Previous Methods
Comparison Methods: We compare the proposed H2T with previous state-of-the-art (SOTA) methods. Two-stage methods, i.e., LDAM-DRW [4], decoupling representation (DR) [19], MisLAS [55] and GCL [27] are compared. DR is the first to decouple the learning of representation and classifier. Zhang et al. [55] and Zhong et al. [56] found that MixUp (MU) [53] can enhance representation learning, therefore we reproduce decoupling representation cooperated with MU for comparison. We also compare with the latest proposed data augmentation methods, i.e., MBJ [31] and CMO [36]. CMO is also an additive approach. We report the results of CE loss and balanced softmax cross-entropy loss (BSCE) [38] with CMO (abbreviated as CE+CMO and BSCE+CMO, respectively). The results of H2T accompanied by DR (CE loss is utilized), MisLAS, and GCL are reported. For multi-expert models, we compare BBN [57], RIDE [47], ACE [3], and ResLT [10]. Both RIDE and ResLT use 3-expert architectures. For RIDE in imageNet-LT and iNaturalist 2018, we report the results of 90 and 180 epochs (eps for short). For Places-LT, we report the results of RIDE with linear classifier and cosine classifier. The results of H2T integrated with RIDE and ResLT are presented for comparison.
Comparison Results: The results on CIFAR10/100-LT are shown in Table 1. H2T can further improve both single and multi-expert models. Applying H2T to the vanilla training of DR with CE loss (DR+H2T) enhances the performance by a significant margin, surpassing most recent methods. The biggest improvement exceeds 5% (on CIRAR-10-LT with $\rho=200$). H2T can also further boost the performance of two-stage SOTA methods, i.e., MisLAS and GCL, albeit with less significant improvements than for DR. H2T also bolsters the efficacy of multi-expert models. For example, RIDE and ResLT integrating with H2T outperform the original methods by 1.01% and 0.87% on CIFAR100-LT with $\rho=100$.
Table 2: Comparison results on imageNet-LT.
Backbone: ResNet-50
Method Year Head Med Tail All
Single Model
CE loss - 64.91 38.10 11.28 44.51
LDAM-DRW [4] 2019 58.63 48.95 30.37 49.96
DR+MU [19] 2020 62.93 49.77 33.26 52.18
MisLAS [56] 2021 62.53 49.82 34.74 52.29
${}^{\star}$ [31] 2022 61.60 48.40 39.00 52.10
GCL [27] 2022 62.24 48.62 52.12 54.51
${}^{\star}$ [36] 2022 67.00 42.30 20.50 49.10
BSCE [${}^{\star}$ [36] 2022 62.00 49.10 36.70 52.30
[…]
MisLAS+H2T - 62.42 51.07 35.36 52.90 ($\uparrow$ 0.61)
GCL+H2T - 62.36 48.75 52.15 54.62 ($\uparrow$ 0.11)
Multi-Expert Model
BBN [57] 2020 - - - 48.30
RIDE [47] (90 eps) 2021 64.33 51.73 36.14 54.08
RIDE [47] (180 eps) 2021 69.59 53.06 30.09 55.72
ACE [${}^{\star}$ 2021 - - - 54.70
ResLT [10] 2022 59.39 50.97 41.29 52.66
[…]
RIDE+H2T (180 eps) - 67.55 54.95 37.08 56.92 ($\uparrow$ 1.20)
ResLT+H2T - 62.29 52.29 35.31 53.39 ($\uparrow$ 0.73)
  • Note: The same as Table 1.
Table 3: Comparison results on iNaturalist 2018.
Backbone: ResNet-50
Method Year Head Med Tail All
Single Model
CE loss - 76.10 69.05 62.44 66.86
LDAM-DRW [4] 2019 - - - 68.15
DR+MU [19] 2020 72.88 71.15 69.24 70.49
MisLAS [56] 2021 72.52 72.08 70.76 71.54
${}^{\star}$ [31] 2022 - - - 70.00
GCL [27] 2022 66.43 71.66 72.47 71.47
${}^{\star}$ [36] 2022 76.90 69.30 66.60 68.90
BSCE [${}^{\star}$ [36] 2022 68.80 70.00 72.30 70.90
[…]
MisLAS+H2T - 69.68 72.49 72.15 72.05 ($\uparrow$ 0.51)
GCL+H2T - 67.74 71.92 72.22 71.62 ($\uparrow$ 0.15)
Multi-Expert Model
BBN [57] 2020 - - - 69.70
RIDE [47] (90 eps) 2021 72.05 73.34 70.23 71.83
RIDE [47] (180 eps) 2021 76.52 74.23 70.45 72.80
ACE [${}^{\star}$ 2021 - - - 72.90
ResLT [${}^{\star}$ 2022 64.85 70.64 72.11 70.69
[…]
RIDE+H2T (180 eps) - 75.69 74.22 71.36 73.11 ($\uparrow$ 0.31)
ResLT with H2T - 68.41 72.31 72.09 71.88 ($\uparrow$ 1.19)
  • Note: The same as Table 1.
Table 4: Comparison results on Places-LT.
Backbone: ResNet-152
Method Year Head Med Tail All
Single Model
CE loss - 46.48 25.66 8.09 29.43
DR+MU [19] 2020 41.66 37.79 32.77 37.40
MisLAS [56] 2021 41.95 41.88 34.65 40.38
${}^{\star}$ [31] 2022 39.50 38.20 35.50 38.10
GCL [27] 2022 38.64 42.59 38.44 40.30
[…]
MisLAS+H2T - 41.40 43.04 35.95 41.03 ($\uparrow$ 0.65)
GCL+H2T - 39.34 42.50 39.46 40.73 ($\uparrow$ 0.43)
Multi-Expert Model
LFME+OLTR [48] ${}^{\star}$ 2020 39.30 39.60 24.20 36.20
RIDE [47] (LC) 2021 44.79 40.69 31.97 40.32
RIDE [47] (CC) 2021 44.38 40.59 32.99 40.35
ResLT [10] 2022 40.30 44.40 34.70 41.00
[…]
RIDE+H2T (CC) - 42.34 43.21 35.62 41.30 ($\uparrow$ 0.95)
  • Note: CC, cosine classifier. LC, linear classifier. Others are the same as Table 1.
We further validate H2T on large-scale datasets, including imageNet-LT, iNaturalist 2018 and Places-LT. The results are summarized in Table 2, Table 3 and Table 4, respectively. The baseline (CE loss) achieves high accuracy on frequent classes while the performance on rare classes is not satisfactory. Compared with the baseline, all methods except CMO and RIDE achieve a significant improvement in the medium and tail classes but at the expense of reduced accuracy on head classes. CMO uses the background of head classes to augment tail classes without reducing the training samples in head classes. CMO accompanying basic CE loss, therefore, improves model performance for both head and tail classes. However, CMO+CE shows less competitive results in improving the classification performance of tail classes. H2T can deliver consistent performance improvements for both single and multi-expert models, particularly for tail and medium classes. Nevertheless, the accuracy improvements of cosine classifier (CC) are less pronounced than those of linear classifier (LC). For example, we can observe that H2T improves GCL that uses CC much less than DR and MisLAS that use LC. In addition, introducing the smooth class margin in GCL can alleviate the sample aliasing between classes in the embedding space. Thus, the enhancement effect of H2T on GCL is less obvious. For example, GCL+H2T improves only by 0.11 on imageNet-LT. Table 4 further compares the improvements of H2T on RIDE with LC (by 1.06%) and CC (0.95%). Note that the margin on LC is more profound.
4.4 Analysis
In this section, We visualize the augmentation effect of H2T on each class, and the experiments are conducted on CIFAR10-LT for a more convenient and clearer presentation. Additionally, we investigate the effects of the sampler used for the fusing branch and the fusion ratio $p$, which are conducted on CIFAR100-LT. All ablations are performed with $\rho$ set to 100, and the DR+H2T method is employed.
Visualization of Feature Space: Figure 4 shows the t-SNE visualization of the distribution in embedding space, which demonstrates our motivation in Section 3.2 that H2T can enrich tail classes so that prevents over-squeezing and overfitting. The features are extracted from class-balanced sampling data. We can see that the distribution of tail classes is sparser than that of head classes. H2T enriches the diversity of intra- and inter-class tail classes with the same number of samples without external information, effectively alleviating the problem of tail class over-squeezing. It is worth noting that, a clear margin, namely the left of Figure 4 is superior in a balanced dataset since the classifier is unbiased for each class. In an imbalanced dataset, a clear margin gives the classifier room to squeeze tail classes in order to correctly classify the head class samples with overwhelming samples. Filling the marge with semantic samples can prevent over-squeezing for tail classes.
Figure 4: Feature distribution comparison w.o. and w. H2T.
The Influence of Sampler for Fusing Branch:
Table 5: Impact of different samplers.
Method Head Med Tail All
CE ${}^{\dagger}$ 67.86 38.51 5.66 39.55
DR w.o. ${}^{\dagger}$ 63.67 46.63 22.21 45.68
[…]
BS+BS 62.72 49.14 25.93 47.30
BS+IS 62.00 50.17 27.07 47.73
  • Note: ${}^{\dagger}$, baselines.
This section validates the efficacy of the fusion strategy proposed in Section 3.3. We compare different sampling strategies including class-balanced sampling (CS), reverse sampling (RS) that is tail-biased, and instance-wise sampling (IS) that is head-biased for the fusing branch. DR uses cRT to fine-tune the classifier with the class-balanced sampling data, which enlarges medium and tail class expands by pushing the decision boundary in the direction of compressing head classes. Additionally, DR uses CE loss that does not consider class margin. These lead to a significant boost in medium and tail classes while reducing the classification accuracy of head classes. RS augments classes of all scales using head class samples with the lowest probability. CS makes all classes augmented with other classes with equal probability. IS allows more features from the medium and tail class samples to be fused with the head features. All sampling strategies adjust the class boundaries towards narrowing the head class distribution. RS has less damage to the performance of head classes while the improvement brought about by IS to the tail and medium classes is more prominent. As the classifier tuning stage cannot change the embedding distribution, the performance on medium and tail classes is further improved but with performance degradation of head classes in the frozen embedding space.
The Impact of Fusion Ratio:
Figure 5: Change of accuracy on different splits w.r.t. $p$.
Figure 5 shows the change of accuracy on different splits with respect to $p$. As the fusion ratio increases, the performance of the medium and tail classes improves while that of the head class decreases, which is consistent with Section 3.4. The class extension of the medium class and the tail class is expanded by eroding that of the head classes through the feature fusion operation. Since $\ddot{w}_{t}^{T}\ddot{f}_{h}>\ddot{w}_{h}^{T}\ddot{f}_{h}$ plays a disproportionate role, resulting in excessive head class samples being incorrectly labeled as medium and tail classes When $p$ increases. As depicted in Figure 5, when $p>0.3$, the accuracy on head classes drops dramatically. Notably, even replacing all feature maps with another class still yields a relatively satisfactory accuracy (43.57%), which surpasses CE loss by 4.02%. $p=0.3$ obtains the best overall performance (47.73%) on CIFAR100-LT. By adjusting $p$ based on specific scenarios, one can control the accuracy of classes with different scales, which is particularly useful when additional accuracy requirements exist for different scale classes. More visualization experiment results can be found in Supplementary A.
5 Concluding Remarks
[…]
Despite its effectiveness, H2T has an underlying assumption that the representation learning has reached its optimum on the available data. It does not alter the feature distribution and only adjusts the decision boundary. Ultimately, improvements in the tail classes are always accompanied by sacrificing the head class performance. Our future work will focus on obtaining a more reasonable embedding space distribution to overcome this limitation.
Supplementary A Visualization Results w.r.t. Different $p$
Figure S1: Visualization of feature distribution w.r.t. $p$.
This supplementary material includes visualization results for fusion ratio $p$ mentioned in Section 3.3 and Section 4.4. We use CIFAR10-LT for easy visualization. Figure S1 shows that as the value of $p$ increases, the interior of each category becomes saturated first, and after $p$ surpasses 0.4, the samples within each category gradually become more dispersed. These findings suggest that the selection of an appropriate fusion ratio is crucial in balancing the preservation of the interior samples while ensuring that the model is able to capture the tail samples effectively. By gaining a deeper understanding of the relationship between the fusion ratio and the distribution of sample points, we can optimize the performance of the model and ensure its suitability for a range of applications.

### v2

@@
-Feature Fusion from Head to Tail: an Extreme Augmenting Strategy for Long-Tailed Visual Recognition
+Feature Fusion from Head to Tail for Long-Tailed Visual Recognition
 Abstract
-The imbalanced distribution of long-tailed data poses a challenge for deep neural networks, as models tend to prioritize correctly classifying head classes over others so that perform poorly on tail classes.
-The lack of semantics for tail classes is one of the key factors contributing to their low recognition accuracy.
-To rectify this issue, we propose to augment tail classes by borrowing the diverse semantic information from head classes, referred to as head-to-tail fusion (H2T).
-We randomly replace a portion of the feature maps of the tail class with those of the head class.
-The fused feature map can effectively enhance the diversity of tail classes by incorporating features from head classes that are relevant to them.
-The proposed method is easy to implement due to its additive fusion module, making it highly compatible with existing long-tail recognition methods for further performance boosting.
+The imbalanced distribution of long-tailed data presents a considerable challenge for deep learning models, as it causes them to prioritize the accurate classification of head classes but largely disregard tail classes.
+The biased decision boundary caused by inadequate semantic information in tail classes is one of the key factors contributing to their low recognition accuracy.
+To rectify this issue, we propose to augment tail classes by grafting the diverse semantic information from head classes, referred to as head-to-tail fusion (H2T).
+We replace a portion of feature maps from tail classes with those belonging to head classes.
+These fused features substantially enhance the diversity of tail classes.
+Both theoretical analysis and practical experimentation demonstrate that H2T can contribute to a more optimized solution for the decision boundary.
+We seamlessly integrate H2T in the classifier adjustment stage, making it a plug-and-play module.
+Its simplicity and ease of implementation allow for smooth integration with existing long-tailed recognition methods, facilitating a further performance boost.
 Extensive experiments on various long-tailed benchmarks demonstrate the effectiveness of the proposed H2T.
-The source code is temporarily available at https://github.com/Keke921/H2T.
+The source code is available at https://github.com/Keke921/H2T.
 1 Introduction
-(a) Comparison of class boundaries. The orange line represents the decision boundary. $\bullet$ and $\blacktriangle$ denotes head and tail class samples, respectively.
-(b) Frequency of predictive labels for tail classes samples. A large number of tail class samples are incorrectly recognized as head classes.
-Figure 1: The relationship between tail and head classes.
-Deep neural networks (DNNs) have shown remarkable capabilities in diverse visual recognition tasks [16, 23, 6, 60, 34], yet their performance heavily relies on data that evenly distribute across categories.
-In contrast, real-world data typically exhibit long-tailed distributions [37, 33, 54], which has become one of the bottlenecks limiting the advancement of DNNs and hinders their broader practical applications.
-For example, misdiagnosis of rare diseases by computer-aided diagnosis [45, 2] may result in irreversible consequences in severe cases.
-Recently, a large number of methods have been proposed to address the class imbalance problem in long-tailed data.
-Class-balancing methods, including re-sampling [5, 18, 39, 20, 35] and re-weighting [17, 12, 4, 30, 15], are the most straightforward approaches to tackle class imbalance.
-Previous class-balancing methods, however, improve the overall performance but are usually at the risk of overfitting to the minority classes or the loss of information [52, 36, 20].
-To overcome these limitations, recent studies have proposed alternative techniques such as re-margining methods [4, 27, 26] and ensembling learning [47, 24, 57, 10].
-These methods can generate more robust predictions, but they do not consider enhancing the sample diversity of tail classes.
-However, insufficient samples and a lack of tail class semantics will also lead to a low recognition rate for tail classes.
-Data augmentation methods [46, 36, 20, 7, 28] are thus proposed to effectively enrich tail classes.
-While these methods have shown state-of-the-art performance, they generally aim to obtain a more representative and separable feature space and do not dig into the potential of the classifier in the obtained feature space.
-This paper thereby aims to unleash the power of the obtained feature space to get a better classifier.
-Take the binary case in the obtained feature space as an example, the semantic space can be divided into three areas: head-like, tail-like, and hard negative.
-The former two areas represent the main distribution spaces of the head and tail classes, respectively, while the hard negative area includes samples that are difficult to distinguish between the two classes.
-In long-tail data, the head class is usually sufficiently sampled, and its semantic area is fully occupied.
-Besides, some samples of head class are distributed in the hard negative area.
-However, only a few samples are often distributed in the tail class semantic regions due to insufficient tail class sampling, resulting in a decision boundary that is limited to the semantic area of the tail class, as shown in the upper of Figure 1a.
-To address the issue of the lack of tail class semantics, we propose a direct and effective solution, named head-to-tail fusion (H2T), which borrows partial of the semantics from the head class and applies them to the tail class.
-Our approach is based on the assumption that predictions on rare instances are easily affected by objects that appear frequently.
-Figure 1b shows a simple experiment [^1] The experiment was conducted on CIFAR10-LT with an imbalance ratio of 100.
-The backbone network is ResNet-32 which was trained using cross-entropy loss.
-Classes 7-9 (0-indexed) are tail classes.for this assumption.
-The label frequency of predicting tail classes into each class is presented.
-We can observe that a large number of tail classes are misclassified as head classes by the model, which may indicate that the hard negative area for tail classes is filled with head class samples.
-Therefore, there is relatively more common closet semantics between head and tail classes.
-By transferring such semantics, we can effectively fill the tail semantic area and improve the recognition rate of the tail class samples, as shown in the lower of Figure 1a.
-On the other hand, by obtaining partial specific semantics (not common semantics) of the head class, the tail class samples can occupy the hard negative semantic region, which forces the classification boundary to move from the tail semantic region to the hard negative semantic region.
-This approach brings the classification boundary closer to the theoretical decision boundary, and improves the classification performance for tail classes.
-To streamline the fusion operation, we design a simple and easy-to-implement strategy that takes full use of the obtained features without retraining the backbone.
-Specifically, we randomly substitute several channels of the feature maps of the class-balanced data with that of the head-biased data.
+Deep models have shown remarkable capabilities in diverse visual recognition tasks (Hu et al.
+2019; Cheung, Li, and Zou 2021; Minaee et al.
+2022; Lan et al.
+2024), yet their performance heavily relies on data that is evenly distributed across categories.
+In contrast, real-world data typically follows long-tailed distributions (Reed 2001; Liu et al.
+2019; Zhang et al.
+2023), which hinders model performance, especially on minority classes (referred to as tail classes in long-tailed data).
+This challenge has become one of the bottlenecks limiting the advancement of deep models.
+In recent times, a plethora of methods have emerged to address the problem of severe class imbalance in long-tailed data from different aspects, such as class-balancing methods (Chawla et al.
+2002; Huang et al.
+2016; Ren et al.
+2018; Cui et al.
+2019; Lin et al.
+2020; Huang et al.
+2020; Hong et al.
+2021; Park et al.
+2021), re-margining methods (Cao et al.
+2019; Li, Cheung, and Lu 2022; Li, Cheung, and Hu 2022), data augmentation (Kim, Jeong, and Shin 2020; Wang et al.
+2021a; Li et al.
+2021; Park et al.
+2022), and ensembling learning (Wang et al.
+2021b; Li et al.
+2022a; Jin et al.
+2023).
+While these methods demonstrated the capability to yield robust predictions, they primarily focus on training a new model to acquire a relatively balanced embedding space and/or assembling multiple diverse networks.
+However, they overlook acquiring a more optimal classifier (determining the decision boundary), which is crucial for fully releasing the potential of the acquired backbone.
+Figure 1: Comparison between decision boundaries produced by (a) existing methods and (b) proposed H2T.
+This paper thereby aims to obtain a more effective classifier with the trained backbone.
+It is widely recognized that increasing class margins (Cao et al.
+2019; Deng et al.
+2019) and/or tightening the intra-class space (Wang et al.
+2018, 2017) for the training set can improve the generalization performance of the model (Liu et al.
+2021; Hu et al.
+2023).
+However, while the clear margin is effective for balanced data, it may still have a biased decision boundary for long-tailed data.
+Take the binary case in the embedding space of the obtained backbone as an example, the decision boundary is usually the midline connecting the two class centroids.
+The head class has sufficiently sampled and its embedding space is fully occupied (Xiao et al.
+2021).
+On the opposite, the tail class suffers from a scarcity of samples, leading to sparsely distributed semantic regions.
+The bias in the tail centroid persists due to head squeeze even if the existence of a clear margin.
+As a result, during the inference stage, a considerable number of samples that differ from the training set emerge, causing the erosion of well-defined margins.
+Consequently, numerous tail class samples are misclassified as head classes, as illustrated in Figure 1(a).
+To enrich the sparse tail class semantics and calibrate the bias in tail classes, we propose a direct and effective solution, named head-to-tail fusion (H2T), which grafts partial semantics from the head class on the tail class.
+In particular, the fusion operation in H2T involves the direct replacement of certain features of tail samples with those of head samples.
+This is based on the assumption that predictions on tail classes that have rare instances are easily affected by head classes that appear frequently.
+Transferring the head semantics can effectively fill the tail semantic area and the category overlap, which compels the decision boundary to shift closer to a more optimal one, as shown in Figure 1(b).
+Ultimately, the generalization performance of the tail classes can be improved.
+To streamline the fusion operation, we design an easy-to-implement strategy that takes full use of the obtained features without retraining the backbone.
+Specifically, we randomly substitute several channels in the feature maps of the class-balanced data with those of the head-biased data.
 This makes the feature maps of tail classes with a high probability of being fused with head classes.
-H2T can be flexibly integrated with other methods, as it does not alter the structure of the backbone as well as any other technical details.
+We apply H2T to the classifier tuning stage, and with just a few lines of code and a few epochs of training, remarkable performance improvements can be achieved.
+In addition, H2T does not alter the structure of the backbone or increase the network parameters.
+This characteristic makes it highly adaptable and compatible with various existing techniques, facilitating its seamless incorporation into diverse existing methods.
 The main contributions of this paper are summarized as follows:
-  • We propose a novel H2T that borrows information from head classes and applies it to augment tail classes without requiring additional data or network parameters, which can enrich tail classes and release the potential of obtained feature space.
-  • We devise a simple fusion strategy that partially substitutes the feature maps of an instance-wise sampled data branch with those of a class-balanced sampled branch.
-The adoption of this approach enables H2T to be implemented with ease and without requiring any modifications to the original model structure.
-  • Considerable performance improvement is achieved in both single and multi-expert models by integrating the H2T module into the original structure.
+  • We propose a novel H2T that borrows information from head classes and applies it to augment tail classes without additional data or network parameters, which can release the potential of the well-trained backbone and obtain better decision boundaries.
+  • We devise a simple fusion strategy that partially substitutes the feature maps of an instance-wise sampling branch with those of a class-balanced sampling branch.
+This approach enables H2T to be implemented with ease and without requiring any modifications to the original backbone structure.
+  • Considerable performance improvement is achieved in both single and multi-expert models by integrating the H2T module into the existing structures.
 Extensive experiments on popular benchmarks validate the efficacy of the proposed method.
 2 Related Work
-This section mainly makes an overview of methods based on data augmentation that is the most relevant regime.
+This section mainly makes an overview of methods based on data augmentation which is the most relevant regime.
 Classical Augmentation: Classical augmentation methods including flip, rotate, crop, padding, and color jittering etc.
-[41, 13] have been widely applied in the process of deep model training.
+(Szegedy et al.
+2015; He et al.
+2016) have been widely applied in deep models.
 On long-tailed data, these methods can improve model robustness and prevent overfitting to specific input patterns to a certain extent.
-Recently, AutoAugment [8], Fast AutoAugment [29], Population Based Augmentation [14] and Randaugment [9], etc.
-have been proposed to automatically determine the best data augmentation strategy for a dataset.
-These methods have been shown to increase classification accuracy for DNNs and demonstrated their efficacy on long-tailed data [38, 11, 52, 25].
-MixUp Based Augmentation: MixUp [53] is an effective data augmentation method that linearly combines a pair of row images and their labels.
+Recently, AutoAugment (Cubuk et al.
+2019), Fast AutoAugment (Lim et al.
+2019), Population based augmentation (Ho et al.
+2019) and Randaugment (Cubuk et al.
+2020), etc.
+have been proposed to determine the best data augmentation strategy for a dataset automatically.
+These methods have been proven to be effective in increasing the classification accuracy for DNNs and also have demonstrated their efficacy on long-tailed data (Ren et al.
+2020; Cui et al.
+2021; Li et al.
+2022b).
+MixUp Based Augmentation: MixUp (Zhang et al.
+2018) is an effective data augmentation method that linearly combines a pair of row images and their labels.
 This method adds complexity control to the uncovered space in the data space through linearly interpolating discrete sample points, which reduces the model generalization error.
-Numerous studies [55, 56, 27] have experimentally proven that MixUp can significantly improve long-tail recognition performance.
-CutMix [50] is a variant of MixUp that augments data by cutting out a patch from the input and pasting it onto another image in the training set.
+Numerous studies (Zhang et al.
+2021; Zhong et al.
+2021; Li, Cheung, and Lu 2022) have experimentally proven that MixUp can significantly improve long-tailed recognition performance.
+CutMix (Yun et al.
+2019), a variant of MixUp, augments data by cutting a patch from one input image and pasting it into another image within the training set.
 The ground truth labels are combined in proportion to the patch areas.
 Park et al.
-[36] proposes to mix the foreground patches from tail classes with the background images from head classes via CutMix, enabling the transfer of context-rich background information from head to tail.
+(Park et al.
+2022) propose mixing the foreground patches from tail classes with the background images from head classes via CutMix, enabling the transfer of context-rich background information from head to tail.
 New Images Based Augmentation: In addition to the aforementioned methods, researchers have proposed alternative approaches that generate new informative samples for tail classes.
-One such method is M2m [20], which translates samples from head classes to the tail using a pre-trained auxiliary classifier.
+One such method is M2m (Kim, Jeong, and Shin 2020), which translates samples from head classes to tail classes using a pretrained auxiliary classifier.
 Wang et al.
-[46] utilize the learned encoded variation information from head classes to synthesize images for tail classes.
+(Wang et al.
+2021a) utilize the learned encoded variation information from head classes to synthesize images for tail classes.
 Zada et al.
-[52] directly exploit pure noise images as tail class samples.
+(Zada, Benou, and Irani 2022) directly exploit pure noise images as tail class samples.
 Zhang et al.
-[55] utilize class activation maps (CAM) [58] to separate image foreground and background, followed by augmenting the foreground through flipping, rotating and scaling, etc.
+(Zhang et al.
+2021) utilize class activation maps (CAM) (Zhou et al.
+2016) to separate image foreground and background, followed by augmenting the foreground by flipping, rotating, and scaling, etc.
 The new samples are then generated by overlaying the augmented foreground on the unchanged background.
 Feature Level Augmentation: Data augmentation can also be performed in feature space.
-For example, Manifold MixUp [44] conducts the linear interpolation on the features of input images, which has been shown to yield better results in long-tail learning [55, 57].
+For example, Manifold MixUp (Verma et al.
+2019) conducts the linear interpolation on the features of input images, which has been shown to yield better results in long-tail learning (Zhang et al.
+2021; Zhou et al.
+2020).
 Chu et al.
-[7] use CAM to decompose features into class-generic and class-specific components.
+(Chu et al.
+2020) use CAM to decompose features into class-generic and class-specific components.
 Tail classes are then augmented by combining their class-specific components with the class-generic components.
-Another promising technique is to transfer knowledge from the head classes to the tail classes.
+Another promising technique is to transfer knowledge from head classes to tail classes.
 For example, Yin et al.
-[49] and Liu et al.
-[32] enrich tail classes with the intra-class variance of head classes to balance the class distributions in feature space.
+(Yin et al.
+2019) and Liu et al.
+(Liu et al.
+2020) enrich tail classes with the intra-class variance of head classes to balance the class distributions in feature space.
 Recently, meta-learning-based augmentation has been leveraged to address class imbalance.
 Liu et al.
-[31] design a meta-embedding that uses a memory bank to enrich tail classes.
-MetaSAug [28] learns class-wise covariance matrices by minimizing the loss on a small balanced validation set and subsequently uses the learned matrices to semantically augment tail classes.
+(Liu, Li, and Sun 2022) design a meta-embedding that uses a memory bank to enrich tail classes.
+MetaSAug (Li et al.
+2021) learns class-wise covariance matrices by minimizing loss on a small balanced validation set and subsequently uses the learned matrices to semantically augment tail classes.
 3 Methodology
@@
 We formally define the basic notations used in this paper before going into detail about our proposed method.
-For dataset: We use $\{x,y\}$ to represent one input image and its corresponding label.
-The total number of classes is denoted by $C$, and the training set includes $N$ samples.
-Thus, we have $y\in\{0,1,\cdots,C-1\}$.
-Suppose class $i$ has $n_{i}$ training samples.
+We use $\{x,y\}$ to represent one input image and its corresponding label.
+The total number of classes is denoted by $C$, thus, we have $y\in\{0,1,\cdots,C-1\}$.
+The training set includes $N$ samples.
+Suppose that class $i$ has $n_{i}$ training samples.
 Then $N=\sum_{i}n_{i}$.
-For simplicity, we suppose $n_{0}\geq n_{1}\geq\cdots\geq n_{C-1}$
-For backbone network: Feed $x$ into the DNN backbone, we can obtain its feature maps before the last pooling layer, denoted as $\mathcal{F}=\left[F_{1},F_{2},\cdots,F_{d}\right]\in\mathbb{R}^{w_{F}\times h_{F}\times d}$, where $w_{F}$ and $h_{F}$ represent the width and height of the feature map, respectively.
+For simplicity, we suppose $n_{0}\geq n_{1}\geq\cdots\geq n_{C-1}$.
+Feed $x$ into the backbone, we can obtain its feature maps before the last pooling layer, denoted as $\mathcal{F}=\left[F_{0},F_{1},\cdots,F_{d-1}\right]\in\mathbb{R}^{w_{F}\times h_{F}\times d}$, where $w_{F}$ and $h_{F}$ represent the width and height of the feature map, respectively.
 $d$ is the dimension of features in embedding space.
 The representation after the last pooling layer is ${f}\in\mathbb{R}^{d}$.
-The weight of the linear classifier is represented as $\textbf{W}=[w_{1},w_{2},\cdots,w_{c}]\in\mathbb{R}^{d\times C}$, where $w_{i}$ represents the classifier weight for class $i$.
+The weight of the linear classifier is represented as $\textbf{W}=[w_{0},w_{1},\cdots,w_{C-1}]\in\mathbb{R}^{d\times C}$, where $w_{i}$ represents the classifier weight for class $i$.
 We use the subscripts $h,m$ and $t$ to indicate head, medium and tail classes, respectively.
@@
 Figure 2: Framework of head-to-tail fusion (H2T).
-We begin by analyzing the compression of the tail distribution caused by imbalances in class sizes.
-The standard training method for DNNs typically involves either empirical risk minimization (ERM) [43, 51] or structural risk minimization (SRM) [22].
-Since the extra item in SRM is a function of model parameters, we can analyze the effect of data on DNNs without loss of generality by focusing solely on ERM.
-To facilitate this analysis, we partition the data by class and use the superscript $i$ to indicate their corresponding labels.
-Formally, ERM can be expressed as:
-$\mathcal{R}_{emp}(\phi)=\frac{1}{N}\sum_{i=1}^{C}\sum_{j=1}^{n_{i}}\mathcal{L}(\phi(x^{i}_{j}),i)$ (1)
-where $\mathcal{L}$ is the loss function.
-$n_{h}\gg n_{t}$ where $h<t$ Thus, misclassifying tail classes have much less risk compared with head classes.
-Generally, a considerable number of tail class samples are mislabeled as head classes.
-The straightforward solution involves increasing the importance or frequencies of tail classes.
-Re-weighting (RW) [4] or re-sampling (RS) [57, 1] can increase the risk of misclassification for tail classes by assigning larger weights or higher sampling rates to them.
-This, in turn, increases the risk of overfitting to tail classes.
-To expand the spatial distribution of tail classes and prevent overfitting, it is crucial to enhance the diversity of tail class samples.
-Unfortunately, obtaining more samples is generally infeasible.
-And Figure 1b shows that the tail classes are more prone to confusing with head classes.
-Typically, there are multiple instances of head and tail classes that are intermixed around the decision boundary.
-We leverage these intermixed instances by directly borrowing semantic information from head classes to augment tail classes.
-We have two purposes: i) increase the risk of misclassifying tail class samples to prevent the model from simply labeling them as head classes; ii) transfer the rich semantic information from head to tail classes, thereby improving the diversity of tail classes.
-3.3 Fusing Head Features to Tail
-Due to the imbalanced distribution of class samples, with the head classes being significantly more numerous than the tail classes, the latter tend to be crowded and underrepresented, leading to poor generalization performance on the test set of tail classes.
-As a large amount of tail class samples are misclassified as head classes, there may exist closet semantics between head and tail classes.
-To exploit this abundant closet semantic information, we fuse the features of head classes to the tail.
-This operation achieves the purpose of enriching the tail classes and expanding their feature space distribution.
+Head classes are significantly more numerous than tail classes in long-tailed data, resulting in a biased classifier, which leads to poor performance on the test set of tail classes.
+The straightforward solution involves increasing the importance or frequencies of tail classes  (Cao et al.
+2019; Zhou et al.
+2020; Ando and Huang 2017).
+They can increase the performance of tail classes, nevertheless, they also entail an elevated risk of overfitting.
+To adjust the decision boundary and prevent overfitting, it is crucial to enhance the diversity of tail class samples (Zhang et al.
+2023; Yang et al.
+2022).
+Unfortunately, directly obtaining more samples is generally infeasible.
+We can consider enriching the tail class by maximizing the utilization of existing features.
+Typically, the misclassified samples are unseen instances from the training set, and these instances tend to be interspersed in the vicinity of the decision boundary.
+This pattern provides us an opportunity to augment the diversity within tail classes and populate class margins with semantically meaningful samples.
+This augmentation process contributes to an optimal decision boundary.
+To achieve this, we simulate the potential unseen samples by directly borrowing semantic information from head classes to augment tail classes.
+By doing so, the rich semantic information can be transferred from well-represented head classes to tail classes.
+3.3 Methodology: Fusing Head Features to Tail
+We fuse the features of head classes to the tail to exploit the abundant closet semantic information.
+This operation can enrich the tail classes and expand their embedding space distribution.
 The fusion process is formulated as:
-~$\tilde{\mathcal{F}}=\mathcal{M}_{p}\otimes\mathcal{F}^{t}+(\mathbf{I}-\mathcal{M}_{p})\otimes\mathcal{F}^{h},$ (2)
-where $\mathcal{M}_{p}$ is the mask combined with identity matrices $\mathbf{I}_{m}\in\mathcal{R}^{w_{F}\times h_{F}}$ and zero matrices $\mathbf{O}_{m}\in\mathcal{R}^{w_{F}\times h_{F}}$.
-The parameter $p$ represents the fusion ratio.
-Specially, the number of $\mathbf{I}_{m}$ and $\mathbf{O}_{m}$ is $d\cdot p$ and $d\cdot(1-p)$, respectively.
-$\tilde{\mathcal{F}}$ is then passed through a pooling layer and classifier to predict the corresponding logits $\mathbf{z}=[z_{1},z_{2},\cdots,z_{C}]$.
-Different loss functions such as CE loss, MisLAS [56], or GCL [27], to name a few, can be adopted.
-The backbone $\phi$ can choose the single model [13] as well as multi-expert model [47, 48].
-We exploit the two-stage method and apply H2T in stage II.
-Input: Training set, fusion ration $p$ ;
-Output: Trained model;
-1 Initialize the model $\phi$ randomly ;
-2 for $iter=1$ to $E_{0}$ do
-3       Sampling batches of data $(x,y)\sim\mathcal{T}^{I}$ from the instance-wise sampling data;
-4       Obtain the feature map $\mathcal{F}=\phi_{\theta}(x)$ ;
-5       Calculate logits $z=\mathbf{W}^{T}\mathbf{f}$ and loss $\mathcal{L}_{1}(x,y)$;
-6       $\phi=\phi-\alpha\nabla_{\phi}\mathcal{L}_{1}((x,y);\phi)$.
-7 end for
-8for $iter=E_{0}+1$ to $E_{2}$ do
-9       Sample batches of data $(x^{B},y^{B})\sim\mathcal{T}^{B}$ and $(x,y)\sim\mathcal{T}^{I}$;
-10       Obtain feature maps $\mathcal{F}^{B}=\phi_{\theta}(x^{B})$ and $\mathcal{F}^{I}=\phi_{\theta}(x)$ ;
-11       Fuse feature maps by Eq. 2 to obtain $\tilde{\mathcal{F}}$ ;
-12       Input $\tilde{\mathcal{F}}$ to pooling layer to obtain $\tilde{\mathbf{f}}$, then calculate the logits by $\tilde{z}=\mathbf{W}^{T}\tilde{\mathbf{f}}$ and the loss by $\mathcal{L}_{2}(x^{B},y^{B})$ ;
-13       Froze the parameters of representation learning $\phi^{r}$, and finetune the classifier parameters $\phi^{c}$: $\phi^{c}=\phi^{c}-\alpha\nabla_{\phi^{c}}\mathcal{L}_{2}((x^{B},y^{B});\phi^{c})$.
-14 end for
+~$\tilde{\mathcal{F}}=\mathcal{M}_{p}\otimes\mathcal{F}_{t}+\overline{\mathcal{M}}_{p}\otimes\mathcal{F}_{h},$ (1)
+where $\mathcal{M}_{p}$ is the mask stacked with multiple 1 matrices $\mathbf{I}_{M}=\mathbb{1}^{w_{F}\times h_{F}}$ and 0 matrices $\mathbf{O}_{M}=\mathbb{0}^{w_{F}\times h_{F}}$.
+The total number of all the 1 and 0 matrices is $d$.
+¯$\overline{\mathcal{M}}{p}$ is the complement of $\mathcal{M}{p}$, that is, the indices of 0 matrices in ¯$\overline{\mathcal{M}}{p}$ correspond to the 1 matrix in $\mathcal{M}{p}$, and vice versa.
+The subscript $p$ of the mask matrices represents the fusion ratio.
+Specially, the number of $\mathbf{I}_{M}$ and $\mathbf{O}_{M}$ is $[d\times p]$ and $d-[d\times p]$, respectively.
+[]delimited-[][$\cdot$][ ⋅ ] means rounding operation.
+$\tilde{\mathcal{F}}$ is then passed through a pooling layer and classifier to predict the corresponding logits $\mathbf{z}=[z_{0},z_{1},\cdots,z_{C-1}]$.
+Different loss functions, such as CE loss, MisLAS (Zhong et al.
+2021), or GCL (Li, Cheung, and Lu 2022), to name a few, can be adopted.
+The backbone $\phi$ can be the single model (He et al.
+2016) as well as the multi-expert model (Wang et al.
+2021b; Xiang, Ding, and Han 2020).
+We exploit the two-stage training (Kang et al.
+2020) and apply H2T in stage II.
 Algorithm 1 H2T
-Visual recognition tasks often involve a vast number of categories.
-Manually screening fused categories can be a tedious and time-consuming process.
-We thereby design a simple strategy.
-We sample two versions of data: 1) class-balanced data $\mathcal{T}^{B}$ used to balance the ERM of each class, which is fed into the fused branch, and 2) instance-wise data $\mathcal{T}^{I}$ has a high probability to obtain head class samples, which is fed into the fusing branch.
-The sampling rates $p^{B}_{i}$ and $p^{I}_{i}$ for $\mathcal{T}^{B}$ and $\mathcal{T}^{I}$ are calculated by
-$p^{B}_{i}=\frac{1}{C\cdot n_{i}},\>p^{I}_{i}=\frac{1}{N}.$ (3)
+0:  Training set, fusion ration $p$
+0:  Trained model
+1:  Initialize the model $\phi$ randomly
+2:  for $iter=1$ to $E_{0}$ do
+3:     Sampling batches of data $(x,y)\sim\mathcal{T}^{I}$ from the instance-wise sampling data
+4:     Obtain the feature map $\mathcal{F}=\phi_{\theta}(x)$
+5:     Calculate logits $z=\mathbf{W}^{T}\mathbf{f}$ and loss $\mathcal{L}_{1}(x,y)$
+6:     $\phi=\phi-\alpha\nabla_{\phi}\mathcal{L}_{1}((x,y);\phi)$.
+7:  end for
+8:  for $iter=E_{0}+1$ to $E_{2}$ do
+9:     Sample batches of data $(x^{B},y^{B})\sim\mathcal{T}^{B}$ and $(x,y)\sim\mathcal{T}^{I}$
+10:     Obtain feature maps $\mathcal{F}^{B}=\phi_{\theta}(x^{B})$ and $\mathcal{F}^{I}=\phi_{\theta}(x)$
+11:     Fuse feature maps by Eq. 1 to obtain $\tilde{\mathcal{F}}$
+12:     Input $\tilde{\mathcal{F}}$ to pooling layer to obtain $\tilde{\mathbf{f}}$, then calculate the logits by $\tilde{z}=\mathbf{W}^{T}\tilde{\mathbf{f}}$ and the loss by $\mathcal{L}_{2}(x^{B},y^{B})$
+13:     Froze the parameters of representation learning $\phi^{r}$, and finetune the classifier parameters $\phi^{c}$: $\phi^{c}=\phi^{c}-\alpha\nabla_{\phi^{c}}\mathcal{L}_{2}((x^{B},y^{B});\phi^{c})$.
+14:  end for
+The selection of features to be fused poses a tedious task during training since visual recognition tasks often encompass a vast number of categories, and it cannot be guaranteed that each minibatch contains the required categories.
+We thereby devise a simple strategy to facilitate an easy-to-execute fusion process.
+This strategy involves sampling two versions of data: 1) class-balanced data $\mathcal{T}^{B}$ used to balance the empirical/structural risk minimization (ERM/SRM) of each class, which is fed into the fused branch, and 2) instance-wise data $\mathcal{T}^{I}$ has a high probability to obtain head class samples, which is fed into the fusing branch.
+The sampling rates $p^{B}_{i}$ and $p^{I}_{i}$ for class $i$ within the sets $\mathcal{T}^{B}$ and $\mathcal{T}^{I}$, respectively, are calculated by
+$p^{B}_{i}=\frac{1}{C},\>p^{I}_{i}=\frac{n_{i}}{N}.$ (2)
 Balanced sampling data ensures that each class is sampled with equal probability $\frac{1}{C}$.
 The fewer the number of samples, the higher the probability of being resampled multiple times.
-The original data is distributed with a long tail, resulting in a head-biased dataset where the head classes have a higher probability of being sampled.
-Next, we can use the feature maps $\mathcal{F}^{B}$ and $\mathcal{F}^{I}$ obtained from $\mathcal{T}^{B}$ and $\mathcal{T}^{I}$ to substitute $\mathcal{F}^{t}$ and $\mathcal{F}^{h}$ in Eq.
+The instance-wise sampling branch deals with the original data distribution, resulting in head-biased data.
+Therefore, the head classes have a higher probability of being sampled.
+Next, we can use the feature maps $\mathcal{F}^{B}$ and $\mathcal{F}^{I}$ obtained from $\mathcal{T}^{B}$ and $\mathcal{T}^{I}$ to substitute $\mathcal{F}_{t}$ and $\mathcal{F}_{h}$ in Eq.
 (1).
 By doing so, the features of the repeatedly sampled tail classes will be fused with the head class features with a higher probability.
-Different from the common practice of linearly combining a pair of inputs and their labels for augmentation, we only use the labels of the fused branch, namely the balance-sampled data, as ground truth to utilize the semantics of head classes to enrich tail classes.
+In contrast to the common practice of linearly combining a pair of inputs and their labels for augmentation, we only use the labels of the fused branch, namely the balance-sampled data, as the ground truth.
+By adopting this method, the semantics of head classes are utilized to enrich tail classes.
 The proposed framework is illustrated in Figure 2 and Algorithm 1.
-The key steps are highlighted in rose-red arrow and text, respectively.
+The key steps are highlighted in rose-red.
 3.4 Rationale Analysis
-We explore the theoretical rationality of the seemingly intuitive operation of H2T in depth in this section.
+Although the proposed H2T appears intuitive and straightforward on the surface, it is built on a foundation of rationale.
+We explore its theoretical rationality in depth in this section.
 For ease of analysis, we assume that the feature maps to be fused are rearranged in order without loss of generality.
-Then, after the pooling layer, the feature can be written as $f_{i}^{T}=[\dot{f}_{i}^{T},\ddot{f}_{i}^{T}]$, where $\dot{f}_{i}$ and $\ddot{f}_{i}$ denote the portions of features that are retained and to be fused, respectively $w_{i}^{T}=[\dot{w}_{i}^{T},\ddot{w}_{i}^{T}]$ is the corresponding classifier weights.
+Then, after the pooling layer, the feature can be written as $f_{i}^{T}=[\dot{f}_{i}^{T},\ddot{f}_{i}^{T}]$, where $\dot{f}_{i}$ and $\ddot{f}_{i}$ denote the portions of features that are retained and to be fused, respectively.
+$w_{i}^{T}=[\dot{w}_{i}^{T},\ddot{w}_{i}^{T}]$ is the corresponding classifier weights.
 On the one hand, we expect $z_{t}>z_{h}$ for tail class, so that
 $\begin{split}w_{t}^{T}f_{t}&>w_{h}^{T}f_{t}\>\>\Rightarrow\\ \dot{w}_{t}^{T}\dot{f}_{t}+\ddot{w}_{t}^{T}\ddot{f}_{t}&>\dot{w}_{h}^{T}\dot{f}_{t}+\ddot{w}_{h}^{T}\ddot{f}_{t}.\end{split}$ (3)
-However, as shown in Figure 1b and analyzed in Section 3.2, numerous tail class samples are incorrectly classified into head classes.
-Therefore, the trained model actually predicts $z_{h}>z_{t}$, namely,
+However, as shown in Figure 3, numerous tail class samples are incorrectly classified into head classes. Therefore, the trained model actually predicts $z_{h}>z_{t}$, namely,
 $\dot{w}_{t}^{T}\dot{f}_{t}+\ddot{w}_{t}^{T}\ddot{f}_{t}<\dot{w}_{h}^{T}\dot{f}_{t}+\ddot{w}_{h}^{T}\ddot{f}_{t}.$ (4)
+Figure 3: Frequency of predictive labels for tail classes samples. A large number of tail class samples are incorrectly recognized as head classes.
+Figure 4: Rationale analysis of H2T.
+Forces 1⃝ and 2⃝ are generated by Eq.
+(9) and Eq.
+(7), respectively.
+1⃝ $>$ 2⃝ makes the tail sample to “pull” closer to $w_{t}$ and “push” further away from $w_{h}$, leading to the adjustment of decision boundary and enlargement of the tail class space.
 After fusing the head to the tail in stage II, the feature of tail class is $\tilde{f}_{t}=[\dot{f}_{t},\ddot{f}_{h}]$, and the corresponding logit is $\tilde{z}_{t}$.
@@
 $\ddot{w}_{h}^{T}(\ddot{f}_{t}-\ddot{f}_{h})>\ddot{w}_{t}^{T}(\ddot{f}_{t}-\ddot{f}_{h}).$ (6)
-We use $\ddot{\theta}_{*}$ to represent the angle between $\ddot{w}_{*}$ and $\ddot{f}_{t}-\ddot{f}_{h}$. Eq. (7) is further simplified to
+We use $\ddot{\theta}_{*}$ (*** is $h$ or $t$) to represent the angle between $\ddot{w}_{*}$ and $\ddot{f}_{t}-\ddot{f}_{h}$. Eq. (6) is further simplified to
 $|\ddot{w}_{h}|\cos\ddot{\theta}_{h}>|\ddot{w}_{t}|\cos\ddot{\theta}_{t}.$ (7)
@@
 Consequently, the classifier increases its performance in the tail class, regardless of the value of $p$.
-Section 4.4 also experimentally validates this analysis.
-Figure 3: Rationale analysis of H2T.
-$\bullet$: $x_{h}$.
-$\blacktriangle$: $x_{t}$.
-$\rightarrow$: $w_{h}$.
-$\rightarrow$: $w_{t}$.
-$\rightarrow$: $f_{t}-f_{h}$.
-Forces 1⃝ and 2⃝ are generated by Eq.
-(10) and Eq.
-(8), respectively.
-1⃝ $>$ 2⃝ makes the tail class sample to“pull" closer to $w_{t}$ and “push" further away from $w_{h}$.
-This leads to the enlargement of the tail class space and a reduction in the priority of head class.
+We also experimentally validate this analysis Section 4.4.
 4 Experiments
-Table 1: Comparison results on CIFAR10/100-LT w.r.t. top-1 accuracy (%).
-Dataset CIFAR10-LT CIFAR100-LT
-Backbone: ResNet-32 Imbalance Ratio
-Method Year 200 100 50 200 100 50
-Single Model
-CE loss - 67.24 70.70 74.81 35.99 39.55 45.40
-LDAM-DRW [4] 2019 73.52 77.03 81.03 38.91 42.04 47.62
-DR+MU [19] 2020 73.06 79.15 84.21 41.73 45.68 50.86
-MisLAS [56] 2021 77.17 82.12 85.55 43.45 46.71 52.31
-${}^{\star}$ [31] 2022 - 81.00 86.60 - 45.80 52.60
-GCL [27] 2022 78.56 82.06 85.36 44.76 48.61 53.55
-${}^{\star}$ [36] 2022 - - - - 43.90 48.30
-BSCE [${}^{\star}$ [36] 2022 - - - - 46.60 51.40
-[…]
-MisLAS+H2T - 78.05 ($\uparrow$ 0.88) 83.23 ($\uparrow$ 1.11) 86.72 ($\uparrow$ 1.17) 43.84 ($\uparrow$ 0.39) 47.62 ($\uparrow$ 0.91) 52.73 ($\uparrow$ 0.42)
-GCL+H2T - 79.42 ($\uparrow$ 0.86) 82.40 ($\uparrow$ 0.34) 85.41 ($\uparrow$ 0.05) 45.24 ($\uparrow$ 0.48) 48.88 ($\uparrow$ 0.27) 53.76 ($\uparrow$ 0.21)
-Multi-Expert Model
-BBN [57] 2020 73.47 79.82 81.18 37.21 42.56 47.02
-RIDE [47] 2021 - - - 45.84 50.37 54.99
-ACE [${}^{\star}$ 2021 - 81.20 84.30 - 49.40 50.70
-ResLT [10] 2022 75.29 80.58 84.07 44.26 48.73 53.81
-[…]
-ResLT+H2T - 77.25 ($\uparrow$ 1.96) 81.77 ($\uparrow$ 1.19) 84.99 ($\uparrow$ 0.92) 46.18 ($\uparrow$ 1.92) 49.60 ($\uparrow$ 0.87) 54.39 ($\uparrow$ 0.58)
-  • Note: The results are obtained by re-implementing the official codes or reproducing with the given parameters, except the methods marked by $\star$ that are quoted from the corresponding papers.
-  • The best and the second-best results are shown in underline bold and bold, respectively.
 4.1 Datasets and Metrics
-Dataset: We evaluated our proposed method on five widely-used benchmarks: CIFAR10/100-LT [4], ImageNet-LT [33], iNaturalist 2018 [42], and Places-LT [33].
-CIFAR10/100-LT are the small-scale datasets that are sub-sampled from their balanced versions, namely, CIFAR10/100 [21].
-We used the imbalance ratios ($\rho=\frac{n_{max}}{n_{min}}$) 200, 100, and 50 [7].
-The original versions of imageNet-2012 [40] and Places 365 [59] are also balanced datasets.
+We evaluate our proposed H2T on four widely-used benchmarks: CIFAR100-LT (Cao et al.
+2019), ImageNet-LT (Liu et al.
+2019), iNaturalist 2018 (Van Horn et al.
+2018), and Places-LT (Liu et al.
+2019).
+CIFAR100-LT is a small-scale dataset that is sub-sampled from the balanced version CIFAR100 (Krizhevsky, Hinton et al.
+2009).
+We used the imbalance ratios ($\rho=\frac{n_{max}}{n_{min}}$) 200, 100, and 50 (Chu et al.
+2020).
+The original versions of imageNet-2012 (Russakovsky et al.
+2015) and Places 365 (Zhou et al.
+2017) are also balanced datasets.
 We use the same settings as Liu et al.
-[33] to obtain the long-tailed versions.
-iNaturalist is a natural dataset collected from all over the world and is heavily imbalanced.
-The 2018 version [42] is utilized in our experiment.
-Metrics: Besides top-1 classification accuracy, following Liu et al.
-[33], the accuracy on three partitions: head ($n_{i}>100$), medium ($20<n_{i}\leq 100$) and tail ($n_{i}\leq 20$) on large-scale datasets are also compared.
+(Liu et al.
+2019) to obtain the long-tailed versions.
+iNaturalist is collected from all over the world and is naturally heavily imbalanced.
+The 2018 version (Van Horn et al.
+2018) is utilized in our experiment.
+We also report the comparison results on CIFAR10-LT (Cao et al.
+2019) in the Appendix.
+As for the metrics, besides top-1 classification accuracy, following Liu et al.
+(Liu et al.
+2019), the accuracy on three partitions: head ($n_{i}>100$), medium ($20<n_{i}\leq 100$) and tail ($n_{i}\leq 20$) on large-scale datasets are also compared.
 4.2 Basic Settings
@@
 Stage II is trained with 10 epochs.
-For CIFAR10/100-LT, we refer to the settings in Cao et al.
-[4] and Zhong et al.
-[56].
-The backbone network is ResNet-32 [13].
-Stage I trains for 200 epochs.The initial learning rate ($l_{r}$) is 0.1 and the batch size is 128.
-$l_{r}$ is decayed at the $160^{th}$ and $180^{th}$ epochs by 0.1.
-For imageNet-LT and iNaturalist 2018, we use the commonly used ResNet-50, while for Places-LT, we utilize the ResNet-152 pre-trained on imageNet.
-For other settings, we mainly follow the original papers of the comparison methods.
+For CIFAR100-LT, we refer to the settings in Cao et al.
+(Cao et al.
+2019) and Zhong et al.
+(Zhong et al.
+2021).
+The backbone network is ResNet-32 (He et al.
+2016).
+Stage I trains for 200 epochs.
+The initial learning rate is 0.1 and is decayed at the $160^{th}$ and $180^{th}$ epochs by 0.1.
+The batch size is 128.
+For imageNet-LT and iNaturalist 2018, we use the commonly used ResNet-50.
+For Places-LT, we utilize the ResNet-152 pre-trained on imageNet.
 We reproduce the prior methods that provide the hyper-parameters for a fair comparison.
-For those methods that do not provide hyperparameters or code, we directly report their results presented in their original paper.
-4.3 Comparisons to Previous Methods
-Comparison Methods: We compare the proposed H2T with previous state-of-the-art (SOTA) methods.
-Two-stage methods, i.e., LDAM-DRW [4], decoupling representation (DR) [19], MisLAS [55] and GCL [27] are compared.
-DR is the first to decouple the learning of representation and classifier.
-Zhang et al.
-[55] and Zhong et al.
-[56] found that MixUp (MU) [53] can enhance representation learning, therefore we reproduce decoupling representation cooperated with MU for comparison.
-We also compare with the latest proposed data augmentation methods, i.e., MBJ [31] and CMO [36].
-CMO is also an additive approach.
-We report the results of CE loss and balanced softmax cross-entropy loss (BSCE) [38] with CMO (abbreviated as CE+CMO and BSCE+CMO, respectively).
+For those methods that do not provide hyperparameters or official codes, we directly report their results presented in the original paper.
+4.3 Comparisons to Existing Methods
+Compared Methods: For single model, we compare the proposed H2T with the following three kinds of methods: two-stage methods, i.e., LDAM-DRW (Cao et al.
+2019), decoupling representation (DR) (Kang et al.
+2020) [^1] We reproduce DR with MixUp for a fair comparison., MisLAS (Zhang et al.
+2021), and GCL (Li, Cheung, and Lu 2022); decision boundary adjustment method, i.e., Adaptive Bias Loss (ABL) (Jin, Lang, and Lei 2023); data augmentation methods, i.e., MBJ (Liu, Li, and Sun 2022), and CMO (Park et al.
+2022).
+We report the results of CE loss and balanced softmax cross-entropy loss (BSCE) (Ren et al.
+2020) with CMO (abbreviated as CE+CMO and BSCE+CMO, respectively).
 The results of H2T accompanied by DR (CE loss is utilized), MisLAS, and GCL are reported.
-For multi-expert models, we compare BBN [57], RIDE [47], ACE [3], and ResLT [10].
+For multi-expert models, we compare BBN (Zhou et al.
+2020), RIDE (Wang et al.
+2021b), ACE (Cai, Wang, and Hwang 2021), and ResLT (Cui et al.
+2023).
 Both RIDE and ResLT use 3-expert architectures.
-For RIDE in imageNet-LT and iNaturalist 2018, we report the results of 90 and 180 epochs (eps for short).
-For Places-LT, we report the results of RIDE with linear classifier and cosine classifier.
-The results of H2T integrated with RIDE and ResLT are presented for comparison.
-Comparison Results: The results on CIFAR10/100-LT are shown in Table 1.
+For Places-LT, we report the results of RIDE with linear and cosine classifiers.
+Table 1: Comparison results on CIFAR100-LT.
+Imbalance Ratio 200 100 50
+Single Model
+CE loss 35.99 39.55 45.40
+LDAM-DRW (Cao et al. 2019) 38.91 42.04 47.62
+DR+MU (Kang et al. 2020) 41.73 45.68 50.86
+MisLAS (Zhong et al. 2021) 43.45 46.71 52.31
+${}^{\star}$ (Liu, Li, and Sun 2022) - 45.80 52.60
+GCL (Li, Cheung, and Lu 2022) 44.76 48.61 53.55
+${}^{\star}$ (Park et al. 2022) - 43.90 48.30
+${}^{\star}$ (Park et al. 2022) - 46.60 51.40
+${}^{\star}$ (Jin, Lang, and Lei 2023) - 46.80 52.10
+[…]
+MisLAS+H2T 43.84 47.62 52.73
+GCL+H2T 45.24 48.88 53.76
+Multi-Expert Model
+BBN (Zhou et al. 2020) 37.21 42.56 47.02
+RIDE (Wang et al. 2021b) 45.84 50.37 54.99
+${}^{\star}$ (Cai, Wang, and Hwang 2021) - 49.40 50.70
+${}^{\star}$ (Park et al. 2022) - 50.00 53.00
+ResLT (Cui et al. 2023) 44.26 48.73 53.81
+[…]
+ResLT+H2T 46.18 49.60 54.39
+  • Note: The backbone is ResNet-32. $\star$ denotes the results quoted from the corresponding papers. The best and the second-best results are shown in underline bold and bold, respectively.
+Table 2: Comparison results on imageNet-LT.
+Method Head Med Tail All
+Single Model
+CE loss 64.91 38.10 11.28 44.51
+LDAM-DRW (Cao et al. 2019) 58.63 48.95 30.37 49.96
+DR (Kang et al. 2020) 62.93 49.77 33.26 52.18
+MisLAS (Zhong et al. 2021) 62.53 49.82 34.74 52.29
+${}^{\star}$ (Liu, Li, and Sun 2022) 61.60 48.40 39.00 52.10
+GCL (Li, Cheung, and Lu 2022) 62.24 48.62 52.12 54.51
+${}^{\star}$ (Park et al. 2022) 67.00 42.30 20.50 49.10
+${}^{\star}$ (Park et al. 2022) 62.00 49.10 36.70 52.30
+${}^{\star}$ (Jin, Lang, and Lei 2023) 62.60 50.30 36.90 53.20
+[…]
+MisLAS+H2T 62.42 51.07 35.36 52.90
+GCL+H2T 62.36 48.75 52.15 54.62
+Multi-Expert Model
+BBN (Zhou et al. 2020) - - - 48.30
+RIDE (Wang et al. 2021b) 69.59 53.06 30.09 55.72
+${}^{\star}$  (Cai, Wang, and Hwang 2021) - - - 54.70
+${}^{\star}$ (Park et al. 2022) 66.40 53.90 35.60 56.20
+ResLT (Cui et al. 2023) 59.39 50.97 41.29 52.66
+[…]
+ResLT+H2T 62.29 52.29 35.31 53.39
+  • Note: The backbone is ResNet-50. Others are the same as Table 1.
+Table 3: Comparison results on iNaturalist 2018.
+Method Head Med Tail All
+Single Model
+CE loss 76.10 69.05 62.44 66.86
+LDAM-DRW (Cao et al. 2019) - - - 68.15
+DR (Kang et al. 2020) 72.88 71.15 69.24 70.49
+MisLAS (Zhong et al. 2021) 72.52 72.08 70.76 71.54
+${}^{\star}$ (Liu, Li, and Sun 2022) - - - 70.00
+GCL (Li, Cheung, and Lu 2022) 66.43 71.66 72.47 71.47
+${}^{\star}$ (Park et al. 2022) 76.90 69.30 66.60 68.90
+${}^{\star}$ (Park et al. 2022) 68.80 70.00 72.30 70.90
+${}^{\star}$ (Jin, Lang, and Lei 2023) - - - 71.60
+[…]
+MisLAS+H2T 69.68 72.49 72.15 72.05
+GCL+H2T 67.74 71.92 72.22 71.62
+Multi-Expert Model
+BBN (Zhou et al. 2020) - - - 69.70
+RIDE (Wang et al. 2021b) 76.52 74.23 70.45 72.80
+ACE (Cai, Wang, and Hwang ${}^{\star}$ 72.90
+RIDE+ ${}^{\star}$ (Park et al. 2022) 68.70 72.60 73.10 72.80
+ResLT (Cui et al. ${}^{\star}$ 64.85 70.64 72.11 70.69
+[…]
+ResLT with H2T 68.41 72.31 72.09 71.88
+  • Note: The backbone is ResNet-50. Others are the same as Table 1.
+Table 4: Comparison results on Places-LT.
+Method Head Med Tail All
+Single Model
+CE loss 46.48 25.66 8.09 29.43
+DR (Kang et al. 2020) 41.66 37.79 32.77 37.40
+MisLAS (Zhong et al. 2021) 41.95 41.88 34.65 40.38
+${}^{\star}$ (Liu, Li, and Sun 2022) 39.50 38.20 35.50 38.10
+GCL (Li, Cheung, and Lu 2022) 38.64 42.59 38.44 40.30
+${}^{\star}$ (Jin, Lang, and Lei 2023) 41.50 40.80 31.40 39.40
+[…]
+MisLAS+H2T 41.40 43.04 35.95 41.03
+GCL+H2T 39.34 42.50 39.46 40.73
+Multi-Expert Model
+${}^{\star}$ (Xiang, Ding, and Han 2020) 39.30 39.60 24.20 36.20
+RIDE (Wang et al. 2021b) (LC) 44.79 40.69 31.97 40.32
+RIDE (Wang et al. 2021b) (CC) 44.38 40.59 32.99 40.35
+${}^{\star}$ (Cui et al. 2023) 40.30 44.40 34.70 41.00
+[…]
+RIDE+H2T (CC) 42.34 43.21 35.62 41.30
+  • Note: CC, cosine classifier. LC, linear classifier. The backbone is ResNet-152. Others are the same as Table 1.
+Comparison Results: The results on CIFAR100-LT are shown in Table 1.
 H2T can further improve both single and multi-expert models.
 Applying H2T to the vanilla training of DR with CE loss (DR+H2T) enhances the performance by a significant margin, surpassing most recent methods.
-The biggest improvement exceeds 5% (on CIRAR-10-LT with $\rho=200$).
+The biggest improvement exceeds 2%.
 H2T can also further boost the performance of two-stage SOTA methods, i.e., MisLAS and GCL, albeit with less significant improvements than for DR.
@@
 For example, RIDE and ResLT integrating with H2T outperform the original methods by 1.01% and 0.87% on CIFAR100-LT with $\rho=100$.
-Table 2: Comparison results on imageNet-LT.
-Backbone: ResNet-50
-Method Year Head Med Tail All
-Single Model
-CE loss - 64.91 38.10 11.28 44.51
-LDAM-DRW [4] 2019 58.63 48.95 30.37 49.96
-DR+MU [19] 2020 62.93 49.77 33.26 52.18
-MisLAS [56] 2021 62.53 49.82 34.74 52.29
-${}^{\star}$ [31] 2022 61.60 48.40 39.00 52.10
-GCL [27] 2022 62.24 48.62 52.12 54.51
-${}^{\star}$ [36] 2022 67.00 42.30 20.50 49.10
-BSCE [${}^{\star}$ [36] 2022 62.00 49.10 36.70 52.30
+The results on large-scale datasets are summarized in Tables 2, 3, and 4.
+The baseline (CE loss) achieves high accuracy on frequent classes while the performance on rare classes is unsatisfactory.
+Comparatively, all methods, except CMO and multi-expert models, exhibit substantial improvements in the medium and tail classes but at the expense of reduced accuracy in head classes.
+CMO leverages the background of head classes to augment tail classes without reducing the number of training samples in head classes, thus enhancing model performance for both head and tail classes.
+However, CMO+CE shows less competitive results in improving tail class performance.
+ABL also adjusts the decision boundary, but its effectiveness is not as pronounced as H2T.
+It is worth noting that all comparison methods necessitate training from scratch.
+H2T can outperform the comparison methods on most datasets by re-fining the classifier with only basic CE loss and several training epochs.
+For example, on iNaturalist 2018, DR+H2T achieves 71.81%, which is the highest among all other comparison methods.
+Furthermore, H2T can deliver consistent performance improvements for both single and multi-expert models, particularly for tail and medium classes.
+Nevertheless, the accuracy improvements of cosine classifier (CC) are less pronounced than those of linear classifier (LC).
+For example, we can observe that H2T improves GCL using CC much less than DR and MisLAS that use LC.
+One reason is that the soft margin in GCL allowing samples to fall in the margin between classes can alleviate the decision boundary bias compared with hard margins such as LDAM (Li et al.
+2023).
+This phenomenon also confirms our motivation.
+Table 4 further compares the improvements of H2T on RIDE with LC (by 1.06%) and CC (0.95%).
+The margin on LC is more profound.
+4.4 Further Analysis
+This section visualizes the decision boundary of H2T on each class and investigates the effects of the sampler for the fusing branch and the fusion ratio $p$ mentioned in Section 3.3.
+All experiments are performed with DR+H2T.
 […]
-MisLAS+H2T - 62.42 51.07 35.36 52.90 ($\uparrow$ 0.61)
-GCL+H2T - 62.36 48.75 52.15 54.62 ($\uparrow$ 0.11)
-Multi-Expert Model
-BBN [57] 2020 - - - 48.30
-RIDE [47] (90 eps) 2021 64.33 51.73 36.14 54.08
-RIDE [47] (180 eps) 2021 69.59 53.06 30.09 55.72
-ACE [${}^{\star}$ 2021 - - - 54.70
-ResLT [10] 2022 59.39 50.97 41.29 52.66
-[…]
-RIDE+H2T (180 eps) - 67.55 54.95 37.08 56.92 ($\uparrow$ 1.20)
-ResLT+H2T - 62.29 52.29 35.31 53.39 ($\uparrow$ 0.73)
-  • Note: The same as Table 1.
-Table 3: Comparison results on iNaturalist 2018.
-Backbone: ResNet-50
-Method Year Head Med Tail All
-Single Model
-CE loss - 76.10 69.05 62.44 66.86
-LDAM-DRW [4] 2019 - - - 68.15
-DR+MU [19] 2020 72.88 71.15 69.24 70.49
-MisLAS [56] 2021 72.52 72.08 70.76 71.54
-${}^{\star}$ [31] 2022 - - - 70.00
-GCL [27] 2022 66.43 71.66 72.47 71.47
-${}^{\star}$ [36] 2022 76.90 69.30 66.60 68.90
-BSCE [${}^{\star}$ [36] 2022 68.80 70.00 72.30 70.90
-[…]
-MisLAS+H2T - 69.68 72.49 72.15 72.05 ($\uparrow$ 0.51)
-GCL+H2T - 67.74 71.92 72.22 71.62 ($\uparrow$ 0.15)
-Multi-Expert Model
-BBN [57] 2020 - - - 69.70
-RIDE [47] (90 eps) 2021 72.05 73.34 70.23 71.83
-RIDE [47] (180 eps) 2021 76.52 74.23 70.45 72.80
-ACE [${}^{\star}$ 2021 - - - 72.90
-ResLT [${}^{\star}$ 2022 64.85 70.64 72.11 70.69
-[…]
-RIDE+H2T (180 eps) - 75.69 74.22 71.36 73.11 ($\uparrow$ 0.31)
-ResLT with H2T - 68.41 72.31 72.09 71.88 ($\uparrow$ 1.19)
-  • Note: The same as Table 1.
-Table 4: Comparison results on Places-LT.
-Backbone: ResNet-152
-Method Year Head Med Tail All
-Single Model
-CE loss - 46.48 25.66 8.09 29.43
-DR+MU [19] 2020 41.66 37.79 32.77 37.40
-MisLAS [56] 2021 41.95 41.88 34.65 40.38
-${}^{\star}$ [31] 2022 39.50 38.20 35.50 38.10
-GCL [27] 2022 38.64 42.59 38.44 40.30
-[…]
-MisLAS+H2T - 41.40 43.04 35.95 41.03 ($\uparrow$ 0.65)
-GCL+H2T - 39.34 42.50 39.46 40.73 ($\uparrow$ 0.43)
-Multi-Expert Model
-LFME+OLTR [48] ${}^{\star}$ 2020 39.30 39.60 24.20 36.20
-RIDE [47] (LC) 2021 44.79 40.69 31.97 40.32
-RIDE [47] (CC) 2021 44.38 40.59 32.99 40.35
-ResLT [10] 2022 40.30 44.40 34.70 41.00
-[…]
-RIDE+H2T (CC) - 42.34 43.21 35.62 41.30 ($\uparrow$ 0.95)
-  • Note: CC, cosine classifier. LC, linear classifier. Others are the same as Table 1.
-We further validate H2T on large-scale datasets, including imageNet-LT, iNaturalist 2018 and Places-LT.
-The results are summarized in Table 2, Table 3 and Table 4, respectively.
-The baseline (CE loss) achieves high accuracy on frequent classes while the performance on rare classes is not satisfactory.
-Compared with the baseline, all methods except CMO and RIDE achieve a significant improvement in the medium and tail classes but at the expense of reduced accuracy on head classes.
-CMO uses the background of head classes to augment tail classes without reducing the training samples in head classes.
-CMO accompanying basic CE loss, therefore, improves model performance for both head and tail classes.
-However, CMO+CE shows less competitive results in improving the classification performance of tail classes.
-H2T can deliver consistent performance improvements for both single and multi-expert models, particularly for tail and medium classes.
-Nevertheless, the accuracy improvements of cosine classifier (CC) are less pronounced than those of linear classifier (LC).
-For example, we can observe that H2T improves GCL that uses CC much less than DR and MisLAS that use LC.
-In addition, introducing the smooth class margin in GCL can alleviate the sample aliasing between classes in the embedding space.
-Thus, the enhancement effect of H2T on GCL is less obvious.
-For example, GCL+H2T improves only by 0.11 on imageNet-LT.
-Table 4 further compares the improvements of H2T on RIDE with LC (by 1.06%) and CC (0.95%).
-Note that the margin on LC is more profound.
-4.4 Analysis
-In this section, We visualize the augmentation effect of H2T on each class, and the experiments are conducted on CIFAR10-LT for a more convenient and clearer presentation.
-Additionally, we investigate the effects of the sampler used for the fusing branch and the fusion ratio $p$, which are conducted on CIFAR100-LT.
-All ablations are performed with $\rho$ set to 100, and the DR+H2T method is employed.
-Visualization of Feature Space: Figure 4 shows the t-SNE visualization of the distribution in embedding space, which demonstrates our motivation in Section 3.2 that H2T can enrich tail classes so that prevents over-squeezing and overfitting.
-The features are extracted from class-balanced sampling data.
-We can see that the distribution of tail classes is sparser than that of head classes.
-H2T enriches the diversity of intra- and inter-class tail classes with the same number of samples without external information, effectively alleviating the problem of tail class over-squeezing.
-It is worth noting that, a clear margin, namely the left of Figure 4 is superior in a balanced dataset since the classifier is unbiased for each class.
-In an imbalanced dataset, a clear margin gives the classifier room to squeeze tail classes in order to correctly classify the head class samples with overwhelming samples.
-Filling the marge with semantic samples can prevent over-squeezing for tail classes.
-Figure 4: Feature distribution comparison w.o. and w. H2T.
-The Influence of Sampler for Fusing Branch:
+Figure 5: Decision boundary Comparison of Class 0 and 8 without H2T v.s. with H2T (namely, DR and DR+H2T).
+Figure 6: Change of accuracy on different splits w.r.t. $p$.
+The Influence of Sampler for Fusing Branch: We compare different sampling strategies for the fusing branch, including class-balanced sampling (CS), reverse sampling (RS) that is tail-biased, and instance-wise sampling (IS) that is head-biased.
+The results are shown in Table 5.
+DR uses cRT to fine-tune the classifier with the class-balanced sampling data, which enlarges the medium and tail class expands by pushing the decision boundary in the direction of compressing head classes.
+Additionally, DR uses CE loss that does not consider class margin.
+These lead to a significant boost in medium and tail classes while reducing the classification accuracy on head classes.
+RS augments classes of all scales using head class samples with the lowest probability.
+CS makes all classes augmented with other classes in an equal probability.
+IS allows more features from the medium and tail classes to be fused with head features.
+All sampling strategies adjust the decision boundaries towards narrowing the head class distribution.
+RS has less damage to the performance of head classes while the improvement brought about by IS to the tail and medium classes is more prominent.
+As the classifier tuning stage cannot change the embedding distribution, the performance on medium and tail classes is further improved but with performance degradation of head classes.
 Table 5: Impact of different samplers.
@@
 BS+IS 62.00 50.17 27.07 47.73
-  • Note: ${}^{\dagger}$, baselines.
-This section validates the efficacy of the fusion strategy proposed in Section 3.3.
-We compare different sampling strategies including class-balanced sampling (CS), reverse sampling (RS) that is tail-biased, and instance-wise sampling (IS) that is head-biased for the fusing branch.
-DR uses cRT to fine-tune the classifier with the class-balanced sampling data, which enlarges medium and tail class expands by pushing the decision boundary in the direction of compressing head classes.
-Additionally, DR uses CE loss that does not consider class margin.
-These lead to a significant boost in medium and tail classes while reducing the classification accuracy of head classes.
-RS augments classes of all scales using head class samples with the lowest probability.
-CS makes all classes augmented with other classes with equal probability.
-IS allows more features from the medium and tail class samples to be fused with the head features.
-All sampling strategies adjust the class boundaries towards narrowing the head class distribution.
-RS has less damage to the performance of head classes while the improvement brought about by IS to the tail and medium classes is more prominent.
-As the classifier tuning stage cannot change the embedding distribution, the performance on medium and tail classes is further improved but with performance degradation of head classes in the frozen embedding space.
-The Impact of Fusion Ratio:
-Figure 5: Change of accuracy on different splits w.r.t. $p$.
-Figure 5 shows the change of accuracy on different splits with respect to $p$.
-As the fusion ratio increases, the performance of the medium and tail classes improves while that of the head class decreases, which is consistent with Section 3.4.
-The class extension of the medium class and the tail class is expanded by eroding that of the head classes through the feature fusion operation.
-Since $\ddot{w}_{t}^{T}\ddot{f}_{h}>\ddot{w}_{h}^{T}\ddot{f}_{h}$ plays a disproportionate role, resulting in excessive head class samples being incorrectly labeled as medium and tail classes When $p$ increases.
-As depicted in Figure 5, when $p>0.3$, the accuracy on head classes drops dramatically.
-Notably, even replacing all feature maps with another class still yields a relatively satisfactory accuracy (43.57%), which surpasses CE loss by 4.02%.
+The Impact of Fusion Ratio: Figure 6 shows the change of accuracy on different splits with respect to $p$.
+As $p$ increases, the performance of medium and tail classes improves while that of head classes decreases, which is consistent with Section 3.4.The class extension of the medium class and the tail class is expanded by eroding that of head classes through H2T fusion.
+Since $\ddot{w}_{t}^{T}\ddot{f}_{h}>\ddot{w}_{h}^{T}\ddot{f}_{h}$ plays a disproportionate role, resulting in excessive head class samples being incorrectly labeled as medium and tail classes as $p$ increases.
+If $p>0.3$, the accuracy on head classes drops dramatically.
+Even replacing all feature maps with another class still yields a relatively satisfactory accuracy (43.57%), surpassing CE loss by 4.02%.
 $p=0.3$ obtains the best overall performance (47.73%) on CIFAR100-LT.
 By adjusting $p$ based on specific scenarios, one can control the accuracy of classes with different scales, which is particularly useful when additional accuracy requirements exist for different scale classes.
-More visualization experiment results can be found in Supplementary A.
+More visualization results for the fused embedding space and the selection strategy of fused features can be found in Section E in the Appendix.
 5 Concluding Remarks
 […]
-Despite its effectiveness, H2T has an underlying assumption that the representation learning has reached its optimum on the available data.
-It does not alter the feature distribution and only adjusts the decision boundary.
+Despite its effectiveness, H2T has an underlying assumption that representation learning has reached its optimum on the available data.
+It only adjusts the decision boundary without altering the feature distribution.
 Ultimately, improvements in the tail classes are always accompanied by sacrificing the head class performance.
 Our future work will focus on obtaining a more reasonable embedding space distribution to overcome this limitation.
-Supplementary A Visualization Results w.r.t. Different $p$
-Figure S1: Visualization of feature distribution w.r.t. $p$.
-This supplementary material includes visualization results for fusion ratio $p$ mentioned in Section 3.3 and Section 4.4.
+Acknowledgments
+This work was supported in parts by NSFC (62306181, U21B2023, 62376233), DEGP Innovation Team (2022KCX TD025), NSFC-RGC (N_HKBU214/21), RGC (12201321, 12202622), and Shenzhen Science and Technology Program (KQTD20210811090044003, RCJC20200714114435012).
+Appendix: Feature Fusion from Head to Tail for Long-Tailed Visual Recognition
+Appendix A Key Execution Code
+H2T is easy to implement and can be seamlessly integrated with various other methods. Below, we outline the key codes of the algorithm:
+Listing 1 PyTorch implementation of H2T
+1def H2T(x1, x2, rho = 0.3):
+2    if type(x1).__name__ == ’Tensor’:
+3        fea_num = x1.shape[1]
+4        index = torch.randperm(fea_num).cuda()
+5        slt_num = int(rho*fea_num)
+6        index = index[:slt_num]
+7        x1[:,index,:,:] = x2[:,index,:,:]
+8    else:
+9        for i in range(len(x1)):
+10            fea_num = x1[i].shape[1]
+11            index = torch.randperm(fea_num).cuda()
+12            slt_num = int(rho*fea_num)
+13            index = index[:slt_num]
+14            x1[i][:,index,:,:] = x2[i][:,index,:,:]
+15    return x1
+Appendix B Comparison Results on CIFAR-10-LT
+In Section 4.3 of the main paper, we show the results on most of the benchmark datasets except CIRAT10-LT.
+Table 6 presents the supplementary results on CIFAR10-LT w.r.t Top-1 classification accuracy (%).
+The improvement of applying H2T to the vanilla training of DR with CE loss (DR+H2T) is more clear.
+The biggest improvement exceeds 5% on CIFAR10-LT with $\rho=200$.
+The improvement in GCL+H2T is less compared with other methods because GCL uses the distance in cosine form.
+This result is consistent with the experimental results on Places-LT (refer to Table 4 of the main paper), namely the H2T improves RIDE with cosine classifier less than with linear classifier.
+Table 6: Comparison results on CIFAR10-LT.
+Imbalance Ratio 200 100 50
+Single Model
+CE loss 67.24 70.70 74.81
+LDAM-DRW (Cao et al. 2019) 73.52 77.03 81.03
+DR+MU (Kang et al. 2020) 73.06 79.15 84.21
+MisLAS (Zhong et al. 2021) 77.17 82.12 85.55
+${}^{\star}$ (Liu, Li, and Sun 2022) - 81.00 86.60
+GCL (Li, Cheung, and Lu 2022) 78.56 82.06 85.36
+[…]
+MisLAS+H2T 78.05 ($\uparrow$ 0.88) 83.23 ($\uparrow$ 1.11) 86.72 ($\uparrow$ 1.17)
+GCL+H2T 79.42 ($\uparrow$ 0.86) 82.40 ($\uparrow$ 0.34) 85.41 ($\uparrow$ 0.05)
+Multi-Expert Model
+BBN (Zhou et al. 2020) 73.47 79.82 81.18
+${}^{\star}$ (Cai, Wang, and Hwang 2021) - 81.20 84.30
+ResLT (Cui et al. 2023) 75.29 80.58 84.07
+[…]
+  • Note: The backbone is ResNet-32.
+$\star$ represents the results quoted from the corresponding papers.
+The best and the second-best results are shown in underline bold and bold, respectively.
+The text in blue indicates the performance improvement compared to the original method w.o.
+H2T.
+Appendix C Selection of fused features
+Table 7: Selection of fused features
+[…]
+47.22 47.23 47.41 47.73
+Random 30%_2 Random %30_3 Random 30%_4 Random 30%_5
+47.75 47.80 47.71 47.66
+The selected fused features are essential.
+We conduct an experiment to show the effect of the selection strategy of fused features.
+The dataset is CIRAR100-LT and the imbalance ratio is 100.
+Table 7 shows the Top-1 classification accuracy (%) of different selections.
+Two approaches are validated: sequential selection and random selection.
+Sequential selection refers to replacing a certain proportion of feature maps directly in accordance with the original order of feature maps.
+We select the first 30%, meddle 30%, and the last 30% of the feature maps to be replaced.
+Random selection refers to randomly selecting a certain proportion of feature maps to replace.
+In the experiment, the fused ratio is 30% and we verified 5 random seeds (Random 30%_1 to Random 30%_5).
+We can observe that random selection generally outperforms sequential selection.
+Different fused features may introduce variations, underscoring the need for a thoughtful selection strategy.
+This aspect remains a topic for our future work.
+Figure 7: Decision boundary Comparison w.o. and w. H2T (all classes).
+Figure 8: Decision boundary Comparison w.o. and w. H2T (Class 0 and 6).
+Figure 9: Decision boundary Comparison w.o. and w. H2T (Class 0 and 7).
+Refer to captionRefer to captionRefer to captionRefer to caption
+Refer to captionRefer to captionRefer to captionRefer to caption
+Refer to captionRefer to captionRefer to captionRefer to caption
+Refer to captionRefer to captionRefer to captionRefer to caption
+Figure 7: Decision boundary Comparison w.o. and w. H2T (all classes).
+Figure 8: Decision boundary Comparison w.o. and w. H2T (Class 0 and 6).
+Figure 9: Decision boundary Comparison w.o. and w. H2T (Class 0 and 7).
+Figure 10: Decision boundary Comparison w.o. and w. H2T (Class 0 and 9).
+Appendix D Visualization Results of decision Boundaries
+This section offers additional visualization comparisons pertaining to the decision boundary, serving as a supplementary analysis to Figure 5 presented in Section 4.4 of the main paper.
+CIFAR10-LT, with an imbalance ratio of 100, is employed for the purpose of facilitating easy visualization.
+Classes 0-3 are head classes and Classes 6-9 are tail classes.
+Figure 10 shows the tSNE of embedding space and the decision boundaries.
+We can see that H2T changes the embedding space of head and tail classes more.
+Figures 10-10 further visualize the decision boundary between head and each tail class.
+Compared with DR, H2T fills the margin between classes and adjusts the decision boundary to a more reasonable position in the obtained embedding space.
+For example, as shown in 10, H2T not only expands the tail class (Class 7) space but also rectifies the head class space (Class 0).
+While this decision boundary enhances the classification effect, it remains suboptimal.
+This presents an intriguing problem worth exploring in future research.
+Appendix E Visualization Results w.r.t. $p$
+This section includes visualization results for fusion ratio $p$ mentioned in Section 3.3 of the main paper.
 We use CIFAR10-LT for easy visualization.
-Figure S1 shows that as the value of $p$ increases, the interior of each category becomes saturated first, and after $p$ surpasses 0.4, the samples within each category gradually become more dispersed.
+Figure 11 shows that as the value of $p$ increases, the interior of each category becomes saturated first, and after $p$ surpasses 0.4, the samples within each category gradually become more dispersed.
 These findings suggest that the selection of an appropriate fusion ratio is crucial in balancing the preservation of the interior samples while ensuring that the model is able to capture the tail samples effectively.
 By gaining a deeper understanding of the relationship between the fusion ratio and the distribution of sample points, we can optimize the performance of the model and ensure its suitability for a range of applications.
+(Figure 11 is located on the next page.)
+Figure 11: Visualization of feature distribution w.r.t. $p$.
+Appendix F Comparison with Existing Methods
+$\bullet$ Clarify H2T.
+H2T constructs virtual tail features to simulate potential sample variations by grafting part of head features into the tail, enabling tail samples to populate the class boundary, as depicted in Fig.
+1.
+It enhances generalization by explicitly introducing perturbations.
+$\bullet$ Comparison with influence-balanced loss (IB) (Park et al.
+2021).
+Reducing decision boundary bias is a common goal in long-tail learning.
+Different from H2T, IB implicitly addresses this issue from an optimization perspective.
+The gradients in stage II of IB is $\lambda_{k}\cdot\left({\nabla\mathcal{L}_{k}}/{\|\nabla\mathcal{L}_{k}\|}\right)$.
+Essentially, IB directly specifies the gradient descent step size to promote a more substantial gradient descent for tail classes to finetune the classifier.
+$\bullet$ Comparison with CMO (Park et al.
+2022).
+CMO allows the transformation of discrete sample values into continuous ones by mixup inputs and labels using CutMix.
+No new samples are created essentially.
+H2T constructs new virtual tail features where labels are not involved in fusion, which can simulate samples that are absent from the training set.
+$\bullet$ Comparison with head-to-tail network (HTTN) (Xiao et al.
+2021).
+HTTN indirectly utilizes head samples to transfer the prototypes.
+It computes the attention ($\alpha_{zj}$) between tail and head prototypes, then uses $\alpha_{zj}$ and the relationship ($W_{transfer}$) between head prototypes and head classifier to identify tail classifier.
+In H2T, tail classes directly employ a portion of head features to simulate unseen tail samples and thus populate the class boundary.
+Appendix G Additional Rational Analysis
+The addition of the inequality obtained from correctly classified tail samples to Eq.
+4 does not eliminate the common terms on both sides of the inequality, and consequently, it does not lead to a meaningful result.
+We consider the inequality relationship obtained by correctly classified tail samples and fused head samples:
+[…]
+Then, we can derive Eq. 9, which encourages $w_{t}$ to be in closer proximity to tail samples, as depicted in Figure 4.